### v1

# Fundamentals of non-parametric statistical inference for integrated quantiles

Abstract. We present a general non-parametric statistical inference theory for integrals of quantiles without assuming any specific sampling design or dependence structure. Technical considerations are accompanied by examples and discussions, including those pertaining to the bias of empirical estimators. To illustrate how the general results can be adapted to specific situations, we derive – at a stroke and under minimal conditions – consistency and asymptotic normality of the empirical tail-value-at-risk, Lorenz and Gini curves at any probability level in the case of the simple random sampling, thus facilitating a comparison of our results with what is already known in the literature. Notes and references concerning underlying technicalities in the case of dependent (i.e., time series) data are offered. As a by-product, our general results provide new and unified proofs of large-sample properties of a number of classical statistical estimators, such as trimmed means, and give additional insights into the origins of, and the reasons for, various necessary and sufficient conditions.
Key words and phrases: integrated quantiles, expected shortfall, tail value at risk, Lorenz curve, Gini curve, trimmed mean, $L$-statistic, distortion risk measure.
1 Introduction
Although well understood, statistical inference for quantiles (a.k.a. values-at-risk) is challenging. This is due to the fact that empirical quantiles are order statistics (e.g., Arnold et al., 2008; David and Nagaraja, 2003), unlike empirical cumulative distribution functions (cdf’s) that are sums of (Bernoulli) random variables and can therefore be easily tackled using classical techniques of mathematical statistics and probability theory (e.g., Shorack, 2017). Inevitably, therefore, in the case of quantiles, quite restrictive assumptions on the population distribution arise. For example, consistency of the empirical quantile requires continuity of the population quantile function at the specified probability level, while asymptotic normality requires conditions on the underlying probability density function (pdf) at the specified quantile (e.g. Serfling, 1980; Shao, 2003).
In many applications that arise in economics, finance, and insurance, quantiles are often integrated. Visualized in Figure 1.1
(a) Actuarial tail-value-at-risk $\mathrm{TVaR}$.
(b) Financial tail-value-at-risk $\mathrm{TV@R}$.
(c) Lorenz curve $\mathrm{LC}$.
(d) Gini curve $\mathrm{GC}$.
Figure 1.1: Curves based on integrated quantiles (details in Section 6).
are four illustrative examples of such integrals and their combinations: they are the actuarial and financial tail-values-at-risk, Lorenz and Gini curves, whose definitions will be given in Section 6. Since, generally speaking, integrals are linear functionals of their integrands, which are quantiles in our case, researchers often use the aforementioned asymptotic results for quantiles and then apply the continuous mapping theorem (e.g., Billingsley, 1999) to derive desired results for integrated quantiles (e.g., Csörgő et al., 1986; Bellini et al., 2022; Han et al., 2024). This approach, however, involves the aforementioned assumptions on the population pdf, although we shall soon see that such assumptions are unnecessary for integrated quantilies; even the very existence of pdf is unnecessary.
Indeed, when developing an asymptotic theory for the Lorenz curve, whose definition is based on integrated quantiles as pointed out by Gastwirth (1971), Csörgő and Zitikis (1996) observed that neither the original formulation of the problem nor the obtained large-sample asymptotic distribution requires the existence of pdf, let alone assumptions on it. Based on the observation, they set out to find a path that would lead to desired asymptotic results without involving pdf’s. They succeeded in achieving this goal by introducing a technical tool that they called the Vervaat process, named after the Dutch mathematician Wim Vervaat, whose pioneering results (Vervaat, 1972a, b) on a combination of the uniform on $[0,1]$ quantiles and empirical processes served an inspiration. We refer to Zitikis (1998) for details and references on the topic.
Note 1.1.
The origin of the research path taken by Csörgő and Zitikis (1996) was humble: it was the equation
$\int_{0}^{1}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u=\int^{\infty}_{-\infty}\big{(}F(x)-G(x)\big{)}\mathrm{d}x$ (1.1)
that is known to hold (see Appendix A.6.1 for details) for every pair of random variables $X$ and $Y$ with finite first moments, where $F$ is the cdf of $X$ (shorthanded as $X\sim F$) and $G$ is the cdf of $Y$ (shorthanded as $Y\sim G$), and $F^{-1}$ and $G^{-1}$ are the corresponding quantile functions. For example, the quantile function $F^{-1}$ of $X$ is given by the equation
$F^{-1}(u)=\inf\{x\in\mathbb{R}\leavevmode\nobreak\ :\leavevmode\nobreak\ F(x)\geq u\}$ (1.2)
for all $u\in(0,1]$, while at the point $u=0$ it is defined as the right-hand limit
$F^{-1}(0)=\lim_{u\downarrow 0}F^{-1}(u).$ (1.3)
(Mathematicians would call $F^{-1}$ the left-continuous generalized inverse of $F$.) Since the quantile function is not the ordinary inverse of $F$, as such may not exist unless the cdf is continuous and strictly increasing, technical difficulties and even overlooks do arise (e.g. Wacker, 2023, and references therein). Note also that $F^{-1}(0)$ and $F^{-1}(1)$ are the two endpoints (finite or infinite) of the support of the cdf $F$, which is the smallest $-\mathbb{R}$ interval that contains all the points $x\in\mathbb{R}$ such that $F(x)\in(0,1)$.
A long series of research articles by various authors followed Csörgő and Zitikis (1996), exploring various theoretical and empirical aspects of integrated quantiles without involving pdf’s. Recently, Wei and Zitikis (2023) suggested an extension of the theory initiated by Csörgő and Zitikis (1996). The current paper takes these developments even further by establishing asymptotic results for integrated quantiles without assuming any specific sampling design or dependence structure between the underlying random variables. In fact, the main results are formulated for generic cdf’s $F$ and them approximating sequences $F_{n}$, $n\in\mathbb{N}$, of cdf’s.
We have organized the rest of the paper as follows. In Section 2, we develop a general statistical inference theory for the integral
$\int_{p}^{1}F^{-1}(u)\mathrm{d}u$ (1.4)
of quantiles over the upper-most layer $(p,1)$ of probabilities, where $p\in(0,1)$ is a fixed boundary probability. We call integral (1.4) the upper-layer integral. It plays a particularly prominent role in insurance, where loss random variables are often non-negative, and thus the right-hand tails of their distributions become of particular concern. Risk measures such as the expected shortfall (ES) and its sister risk measure called the “actuarial” tail-value-at-risk (TVaR) arise (e.g., Denuit et al., 2005). Its formula with an accompanying discussion will be provided in Example 6.1 later in this paper. In Section 3 we shall offer several simulated experiments that will clarify some of the statistical properties of the empirical upper-layer integral.
Before we delve into the topics of the following sections, we next present several notes that are useful for understanding and appreciating the results of this paper, and in particular the assumptions under which the results have been established.
Note 1.2.
When $p=0$, integral (1.4) reduces to the mean
$\mu=\int_{0}^{1}F^{-1}(u)\mathrm{d}u$
of the cdf $F$. Statistical inference for the mean under various sampling designs can, of course, be developed without invoking the existence of the pdf of $F$. When $p=1$, integral (1.4) vanishes. Hence, by restricting ourselves to $p\in(0,1)$ we are just excluding the two statistically trivial cases $p=0$ and $p=1$.
Note 1.3.
There is a deeper reason for separating the boundary case $p=0$ from $p\in(0,1)$. First, when $p=0$, in the case of the simple random sampling (SRS), that is, when we work with $n$ independent and identically distributed (iid) random variables, the empirical estimator of $\mu$ is an asymptotically normal estimator under the only requirement that the second moment of the underlying random variable is finite. No assumption related to continuity of the cdf $F$ at any point of the real line is needed. In the case $p\in(0,1)$, however,
for an empirical estimator of the upper-layer integral $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$ to be asymptotically normal, it is necessary to assume that the quantile function $F^{-1}$ is continuous at the point $p$,
as demonstrated by Stigler (1973) when the empirical estimator is a trimmed mean based on SRS. This assumption is also needed for our general results as they absorb the SRS-based trimmed mean as a special case. No such condition, however, is required for consistency of the empirical estimator of the upper-layer integral $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$. We shall illustrate this phenomenon with a simulated example in Section 3.
Note 1.4.
The aforementioned continuity of the quantile function $F^{-1}$ at the point $p$ permeates the entire paper. In our minds, however, we rarely visualize distributions in terms of quantiles – we often think of them in terms of cdf’s. For this reason it is beneficial to follow Shao (2003) and reformulate the continuity of $F^{-1}$ at the point $p\in(0,1)$ in the form of the bounds
$F\left(F^{-1}(p)-\varepsilon\right)<p<F\left(F^{-1}(p)+\varepsilon\right)$ (1.5)
that need to be satisfied for every $\varepsilon>0$. Figure 1.2
(a) Cdf: The only $p\in(0,1)$ that fails condition (1.5) is $p=p_{2}$, provided that $x_{1}<x_{2}$. When $x_{1}=x_{2}$ (no gap), every $p\in(0,1)$ satisfies the condition.
(b) Quantile function: The only discontinuity of $F^{-1}$ is at $p=p_{2}$, provided that $x_{1}<x_{2}$. When $x_{1}=x_{2}$ (no jump), the quantile function is continuous.
Figure 1.2: An illustrative cdf and its quantile function with the boundary points $p_{1}\leq p_{2}$ and $x_{1}\leq x_{2}$ of the distributional jump and gap, respectively, where $x_{1}=F^{-1}(p_{2})$.
illustrates the reformulation. Note in the figure that $F$ is strictly increasing at the point $F^{-1}(p)$ for every $p\in(0,1)$ except when $p=p_{2}$, in which case there is a flat region adjacent to $x_{1}=F^{-1}(p_{2})$ and, therefore, the quantile function has a jump of size $x_{2}-x_{1}$ at the point $p_{2}$. Given this explanation, for the sake of brevity, throughout the rest of the paper we follow Stigler (1973) and formulate the condition simply as continuity of the quantile function $F^{-1}$ at the point $p$.
In Section 4, we develop a general statistical inference theory for the integral
$\int_{0}^{p}F^{-1}(u)\mathrm{d}u$ (1.6)
of quantiles over the lower-most layer $(0,p)$ of probabilities, where $p\in(0,1)$ is a fixed boundary probability. Throughout the current paper we call integral (1.6) the lower-layer integral, although in the literature the integral is often called the generalized (or absolute) Lorenz curve (e.g., Davydov and Zitikis, 2002; Helmers and Zitikis, 2005; Davydov et al., 2007, and references therein) to reflect the fact that by dividing the integral by the mean $\mu$ of $F$, we obtain the Lorenz curve (see Example 6.3) that permeates the literature on economic inequality. Integral (1.6) also plays a significant role in finance, because financial losses are usually modelled using negative random variables. Risk measures such as the “financial” tail-value-at-risk (TV@R) arise (e.g., Pflug and Römisch, 2007). Its formula with an accompanying discussion will be provided in Example 6.2.
Note 1.5.
Analogous observations to those made in Notes 1.2 and 1.3 apply to integral (1.6) as well: first, we exclude the statistically trivial cases $p=0$ and $p=1$, and second, asymptotic normality of the empirical estimator when $p\in(0,1)$, unlike in the case $p=1$ that gives rise to the mean $\mu$, requires the quantile function $F^{-1}$ to be continuous at the point $p$, as shown by Stigler (1973) in the case of trimmed mean arising from SRS. No such condition is required for consistency of the estimator of the lower-layer integral. Details are in Section 4.
There are also problems, like those associated with insurance layers (e.g, Wang, 1996), when aggregating middle quantiles is of interest. Hence, in Section 5 we develop a general statistical inference theory for the integral
$\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u$ (1.7)
of quantiles over the middle layer $(p_{1},p_{2})$ of probabilities, where $10<p_{1}<p_{2}<$ 1 are fixed boundary probabilities. We call integral (1.7) the middle-layer integral.
Note 1.6.
When $p_{1}=0$ and $p_{2}=1$, integral (1.7) reduces to the mean $\mu$ of the cdf $F$. When $p_{1}=0<p_{2}<1$, integral (1.7) reduces to integral (1.6), and when $10<p_{1}<p_{2}=$ 1, integral (1.7) reduces to integral (1.4). Hence, by restricting our study of the middle-layer integral to the case $10<p_{1}<p_{2}<$ 1 only, we are not losing anything from the statistical point of view.
Note 1.7.
Asymptotic normality of the empirical estimator when $10<p_{1}<p_{2}<$ 1 requires the quantile function $F^{-1}$ to be continuous at the points $p_{1}$ and $p_{2}$, as follows from Stigler (1973) in the case of trimmed mean arising from SRS. No such condition is required for consistency of the estimator. Details are in Section 5.
When modeling both profits and losses (so-called P&L variables), the underlying population distribution may span the entire real line $\mathbb{R}$, and thus both upper- and lower-layer integrals (1.4) and (1.6) may show up simultaneously. There are also problems that give rise to various combinations of the three integrals introduced above, and one of such instances is the Gini curve (GC), whose formula with an accompanying discussion will be provided in Example 6.4. The general results that we shall develop in the following sections are readily applicable to such combinations as well.
As we have already alluded to, in Section 6 we shall illustrate asymptotic results that can be obtained for the actuarial and financial tail-values-at-risk, as well as for the Lorenz and Gini curves. To maximally simplify the illustrations, yet to maintain their informative nature, we shall use SRS, that is, we shall work with iid random variables. We stress, however, that this choice of the sampling design is only to facilitate illustrations of our general results, which do not rely on any particular sampling design or dependence structure.
Section 7 connects the results of previous sections with research areas such as $L$-statistics, distortion and spectral risk measures. Section 8 concludes the paper with a brief summary. Longer proofs, accompanying lemmas, and some cursory technicalities are in Appendix A.
2 The upper-layer integral
In this section we develop statistical inference for the upper-layer integral $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$ at any fixed lower probability level
$p\in(0,1).$
Of course, we need to assume that the cdf $F$ is from the class $\mathcal{L}^{+}_{1}$ of those cdf’s for which the integral $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$ is finite, which is the same as to require that the positive part $X^{+}:=X\vee 0$ of the random variable $X\sim F$ has a finite first moment $\mathbb{E}(X^{+})<\infty$.
The functional
$\mathcal{L}^{+}_{1}\ni F\mapsto\int_{p}^{1}F^{-1}(u)\mathrm{d}u\in\mathbb{R}$
is not linear, which makes it challenging to derive statistical inference. It appears, however, that when developing statistical inference, the functional can be approximated by the linear one
$\mathcal{L}^{+}_{1}\ni G\mapsto\int^{\infty}_{F^{-1}(p)}\big{(}1-G(x)\big{)}\mathrm{d}x\in\mathbb{R}.$ (2.1)
Indeed, we shall see from the following corollaries that the difference between the empirical upper-layer integral $\int_{p}^{1}F_{n}^{-1}(u)\mathrm{d}u$ and its population counterpart $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$ gets asymptotically close to the difference between $\int^{\infty}_{F^{-1}(p)}\big{(}1-F_{n}(x)\big{)}\mathrm{d}x$ and $\int^{\infty}_{F^{-1}(p)}\big{(}1-F(x)\big{)}\mathrm{d}x$, where $F_{n}\in\mathcal{L}^{+}_{1}$, $n\in\mathbb{N}$, are cdf’s approaching $F$ when the parameter $n$ grows indefinitely. The corollaries rely on the following fundamental theorem, whose proof is given in Appendix A.2.
Theorem 2.1.
Let $F$ and $G$ be any two cdf’s from $\mathcal{L}^{+}_{1}$. Then
$\int_{p}^{1}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u=\int^{\infty}_{F^{-1}(p)}\big{(}F(x)-G(x)\big{)}\mathrm{d}x-\mathrm{Rem}(p;F,G),$ (2.2)
where the remainder term
$\mathrm{Rem}(p;F,G)=\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}G(x)-p\big{)}\mathrm{d}x$ (2.3)
is non-negative, that is, $\mathrm{Rem}(p;F,G)\geq 0$, and satisfies the bounds
$\mathrm{Rem}(p;F,G)$ $\leq\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}G(x)-F(x)\big{)}\mathrm{d}x$ (2.4)
$\leq\big{|}G^{-1}(p)-F^{-1}(p)\big{|}\leavevmode\nobreak\ \sup_{x\in\mathbb{R}}\big{|}G(x)-F(x)\big{|}.$ (2.5)
Note 2.1.
The functional $(F,G)\mapsto\mathrm{Rem}(p;F,G)\in[0,\infty)$ is not symmetric. Note that its symmetrization $\mathrm{Rem}(p;F,G)+\mathrm{Rem}(p;G,F)$ is equal to the right-hand side of bound (2.4), that is,
$\mathrm{Rem}(p;F,G)+\mathrm{Rem}(p;G,F)=\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}G(x)-F(x)\big{)}\mathrm{d}x.$ (2.6)
This equation implies bound (2.4) because $\mathrm{Rem}(p;G,F)$ is non-negative, just like $\mathrm{Rem}(p;F,G)$ is.
We are now ready to formulate and discuss several corollaries to Theorem 2.1.
Corollary 2.1 (consistency).
Suppose that $F\in\mathcal{L}^{+}_{1}$, and let $F_{1},F_{2},\ldots\in\mathcal{L}^{+}_{1}$ be any sequence of cdf’s such that
$\sup_{x\in\mathbb{R}}\big{|}F_{n}(x)-F(x)\big{|}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}0$ (2.7)
when $n\to\infty$. Then
$\int_{p}^{1}F_{n}^{-1}(u)\mathrm{d}u-\int_{p}^{1}F^{-1}(u)\mathrm{d}u=\int^{\infty}_{F^{-1}(p)}\big{(}F(x)-F_{n}(x)\big{)}\mathrm{d}x+o_{\mathbb{P}}(1)$ (2.8)
and, therefore, the consistency statement
$\int_{p}^{1}F_{n}^{-1}(u)\mathrm{d}u\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int_{p}^{1}F^{-1}(u)\mathrm{d}u$ (2.9)
holds if and only if
$\int^{\infty}_{F^{-1}(p)}\big{(}1-F_{n}(x)\big{)}\mathrm{d}x\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int^{\infty}_{F^{-1}(p)}\big{(}1-F(x)\big{)}\mathrm{d}x.$ (2.10)
Proof.
The corollary follows from Theorem 2.1 with $F_{n}$ instead of $G$, because $F^{-1}(p)$ is finite and $F_{n}^{-1}(p)$ is asymptotically bounded. Hence, condition (2.7) implies $\mathrm{Rem}(p;F,G)=o_{\mathbb{P}}(1)$ due to bound (2.5). This establishes statement (2.8), which implies the equivalence of statements (2.9) and (2.10), and concludes the proof of Corollary 2.1. ∎
Example 2.1 (SRS).
Let $X_{1},\dots,X_{n}$ be iid random variables. The empirical cdf $F_{n,\textsc{srs}}$ is defined by the equation
$F_{n,\textsc{srs}}(x)={1\over n}\sum_{i=1}^{n}\mathds{1}\{X_{i}\leq x\},$
which is the arithmetic mean of $n$ independent copies of the Bernoulli random variable $\mathds{1}\{X\leq x\}$ taking value $1$ with probability $F(x)$ and 00 otherwise. In other words, $\mathds{1}\{X\leq x\}$ is the indicator that takes value $1$ when the statement $X\leq x$ holds and 00 otherwise. Analogously to equation (1.2), the empirical quantile function $F_{n,\textsc{srs}}^{-1}$ is defined by
$F_{n,\textsc{srs}}^{-1}(u)$ $=\inf\{x\in\mathbb{R}\leavevmode\nobreak\ :\leavevmode\nobreak\ F_{n,\textsc{srs}}(x)\geq u\}$ (2.11)
$=X_{\lceil nu\rceil:n}$ (2.12)
for all $u\in(0,1]$, where $X_{1:n}\leq\cdots\leq X_{n:n}$ are the order statistics of the random variables $X_{1},\dots,X_{n}$, and, for any real $x\in\mathbb{R}$, $\lceil x\rceil$ denotes the smallest integer $k$ such that $x\leq k$, that is, $x\mapsto\lceil x\rceil$ is the classical ceiling function. When $u=0$, we define
$F_{n,\textsc{srs}}^{-1}(0)=X_{1:n}=\min_{1\leq i\leq n}X_{i}.$
Condition (2.7) with $F_{n,\textsc{srs}}$ in place of $F_{n}$ is satisfied by the Glivenko-Cantelli theorem. Hence, we are left to verify condition (2.10). By linearity of functional (2.1), the integral $\int^{\infty}_{F^{-1}(p)}\big{(}1-F_{n,\textsc{srs}}(x)\big{)}\mathrm{d}x$ is the arithmetic mean $n^{-1}\sum_{i=1}^{n}Y_{i,\textsc{upper}}(p)$ of $n$ independent copies of the random variable
$Y_{\textsc{upper}}(p)$ $=\int^{\infty}_{F^{-1}(p)}\mathds{1}\{X>x\}\mathrm{d}x$
$=\big{(}X-F^{-1}(p)\big{)}^{+}.$ (2.13)
The random variable $Y_{\textsc{upper}}(p)$ has a finite first moment because $F\in\mathcal{L}^{+}_{1}$ (see Lemma A.1 for details). Hence, statement (2.10) holds and so, by Corollary 2.1, we have that $\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ is a consistent estimator of $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$. In summary, we conclude that when $F\in\mathcal{L}^{+}_{1}$, we have
$\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int_{p}^{1}F^{-1}(u)\mathrm{d}u.$ (2.14)
Note 2.2.
In addition to equation (2.12), there are also other ways to express $F_{n,\textsc{srs}}^{-1}(u)$ in terms of the order statistics $X_{1:n}\leq\cdots\leq X_{n:n}$. For example, using the integer part $[x]$ of $x\in\mathbb{R}$, in the section on sample quantiles Shao (2003) gives the equation
$F_{n,\textsc{srs}}^{-1}(u)=c_{nu}X_{[nu]:n}+(1-c_{nu})X_{([nu]+1):n},$
where $c_{nu}=1$ if $nu$ is an integer and $c_{nu}=0$ otherwise.
Corollary 2.2 (bias).
Suppose that $F\in\mathcal{L}^{+}_{1}$, and let $F_{1},F_{2},\ldots\in\mathcal{L}^{+}_{1}$ be any sequence of cdf’s that are unbiased estimators of $F$, that is,
$\mathbb{E}\big{(}F_{n}(x)\big{)}=F(x)$ (2.15)
for all $x\in\mathbb{R}$ and $n\in\mathbb{N}$, and such that
$\mathbb{E}\big{(}F_{n}^{-1}(p)\big{)}\in\mathbb{R}$ (2.16)
for all $n\in\mathbb{N}$. Then $\int_{p}^{1}F_{n}^{-1}(u)\mathrm{d}u$ is a non-positively biased estimator of $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$.
Proof.
Using Theorem 2.1 with $F_{n}$ instead of $G$, we have
$\mathrm{Bias}_{n}^{\textsc{upper}}(p):=$ $\mathbb{E}\bigg{(}\int_{p}^{1}F_{n}^{-1}(u)\mathrm{d}u\bigg{)}-\int_{p}^{1}F^{-1}(u)\mathrm{d}u$
$=$ $-\mathbb{E}\big{(}\mathrm{Rem}(p;F,F_{n})\big{)}\in(-\infty,0]$
for every $n\in\mathbb{N}$. Note that $\mathbb{E}\big{(}\mathrm{Rem}(p;F,F_{n})\big{)}$ is finite because of bound (2.5) and condition (2.16). This establishes Corollary 2.2. ∎
Example 2.2 (SRS).
When $F_{n}$ is $F_{n,\textsc{srs}}$, condition (2.15) is obviously satisfied. Condition (2.16) is satisfied when (Stigler, 1974, Proposition 2, p. 679)
there is $\varepsilon>0$ such that $x^{\varepsilon}(1-F(x)+F(-x))\to 0$ when $x\to\infty$, which is equivalent to saying that $x^{\varepsilon}\mathbb{P}(|X|\geq x)\to 0$, where $X\sim F$.
For a related study that connects condition (2.16) with the order of finite moments of $F$, we refer to Gribkova (1995) (see also next Note 2.3 on the topic). Denote the class of all cdf’s satisfying the aforementioned tail-based condition of Stigler (1974) by $\mathcal{T}_{\varepsilon}$. Hence, Corollary 2.2 tells us that when $F\in\mathcal{L}^{+}_{1}\cap\mathcal{T}_{\varepsilon}$ for some $\varepsilon>0$, the estimator $\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ of the upper-layer integral $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$, although being consistent by Example 2.1, has non-positive bias
$\mathrm{Bias}_{n,\textsc{srs}}^{\textsc{upper}}(p):=-\mathbb{E}\big{(}\mathrm{Rem}(p;F,F_{n,\textsc{srs}})\big{)}\in(-\infty,0].$ (2.17)
Note 2.3.
The condition that $F\in\mathcal{L}^{+}_{1}\cap\mathcal{T}_{\varepsilon}$ for some $\varepsilon>0$ can equivalently be reformulated as $F\in\mathcal{L}^{+}_{1}\cap\mathcal{T}^{-}_{\varepsilon}$ for some $\varepsilon>0$, where $\mathcal{T}^{-}_{\varepsilon}$ consists of all those cdf’s that satisfy $|x|^{\varepsilon}F(x)\to 0$ when $x\to-\infty$. This reformulation is possible due to the fact that $F\in\mathcal{L}^{+}_{1}$ implies $x^{\varepsilon}(1-F(x))\to 0$ when $x\to\infty$ for every $\varepsilon\in(0,1)$, and we need this condition to hold for just one $\varepsilon>0$. For those researchers who think in terms of moments, a sufficient condition that ensures the existence of $\varepsilon>0$ such that $F\in\mathcal{L}^{+}_{1}\cap\mathcal{T}^{-}_{\varepsilon}$ would be the requirement that
$\mathbb{E}(X^{+})<\infty\quad\text{and}\quad\mathbb{E}((X^{-})^{\varepsilon})<\infty$
for some $\varepsilon>0$, where $X^{-}=(-X)\vee 0$ is the negative part of the random variable $X\sim F$. Note that this requirement can equivalently be rewritten as $F\in\mathcal{L}^{+}_{1}\cap\mathcal{L}^{-}_{\varepsilon}$.
The next corollary plays a pivotal role in deriving asymptotic distribution of the empirical upper-layer integral. We note at the outset that the corollary crucially relies on the rate of convergence to 00 – denoted by $O_{\mathbb{P}}(1/A_{n})$ – of the supremum in condition (2.7). For example, we would typically have $A_{n}=\sqrt{n}$ for i.i.d. sequences and, more broadly, for weakly-dependent stationary sequences $(X_{n})$, and we would have $A_{n}=n^{\theta}$ with some $\theta\in(0,1/2)$ for long-memory random sequences, as for such sequences, the empirical cdf $F_{n}$ converges to $F$ at a slower rate when $n$ increases. We shall give a more detailed note on the topic near the end of this section. At the moment we emphasize that next Corollary 2.3 does not depend on any sampling design or dependence structure, just like other main results of this paper.
Corollary 2.3 (asymptotic distribution).
Suppose that $F\in\mathcal{L}^{+}_{1}$, and let $F_{1},F_{2},\ldots\in\mathcal{L}^{+}_{1}$ be any sequence of cdf’s such that
$F_{n}^{-1}(p)\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}F^{-1}(p)$ (2.18)
and
$A_{n}\sup_{x\in\mathbb{R}}\big{|}F_{n}(x)-F(x)\big{|}=O_{\mathbb{P}}(1)$ (2.19)
when $n\to\infty$, where $A_{n}\to\infty$ are some normalizing constants. Then
$A_{n}\left(\int_{p}^{1}F_{n}^{-1}(u)\mathrm{d}u-\int_{p}^{1}F^{-1}(u)\mathrm{d}u\right)=A_{n}\int^{\infty}_{F^{-1}(p)}\big{(}F(x)-F_{n}(x)\big{)}\mathrm{d}x+o_{\mathbb{P}}(1)$ (2.20)
and, therefore, the convergence-in-distribution statement
$A_{n}\left(\int_{p}^{1}F_{n}^{-1}(u)\mathrm{d}u-\int_{p}^{1}F^{-1}(u)\mathrm{d}u\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}_{\textsc{upper}}(p)$ (2.21)
holds if and only if
$A_{n}\left(\int^{\infty}_{F^{-1}(p)}\big{(}1-F_{n}(x)\big{)}\mathrm{d}x-\int^{\infty}_{F^{-1}(p)}\big{(}1-F(x)\big{)}\mathrm{d}x\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}_{\textsc{upper}}(p),$ (2.22)
where $\mathcal{L}_{\textsc{upper}}(p)$ is a random variable determined by statement (2.22).
Proof.
Using Theorem 2.1 with $F_{n}$ instead of $G$, we only need to show that $A_{n}\,\mathrm{Rem}(p;F,F_{n})$ converges in probability to 00. Since $\mathrm{Rem}(p;F,F_{n})$ is non-negative and does not exceed
$\big{|}F_{n}^{-1}(p)-F^{-1}(p)\big{|}\leavevmode\nobreak\ \sup_{x\in\mathbb{R}}\big{|}F_{n}(x)-F(x)\big{|},$
the statement $A_{n}\mathrm{Rem}(p;F,F_{n})=o_{\mathbb{P}}(1)$ follows from conditions (2.18) and (2.19). This establishes equation (2.20), which implies the equivalence of statements (2.21) and (2.22), and concludes the proof of Corollary 2.3. ∎
Note 2.4.
The limiting random variable $\mathcal{L}_{\textsc{upper}}(p)$ may or may not be normal. The nature of this random variable is determined by the sequence $(X_{n})$: if it is i.i.d. or, more broadly, weakly dependent and stationary, then $\mathcal{L}_{\textsc{upper}}(p)$ is normal, but if the stationary sequence $(X_{n})$ is long-range dependent, then the limiting random variable $\mathcal{L}_{\textsc{upper}}(p)$, if it exists, is typically non-normal.
Concerning condition (2.18), it is a simple probability exercise to check the validity of the following lemma by imitating related considerations in the section on SRS-based quantiles by Shao (2003).
Lemma 2.1.
When the quantile function $F^{-1}$ is continuous at the point $p$, condition (2.7) implies condition (2.18) and, therefore, the latter condition can be dropped from Corollary 2.3 because condition (2.19) implies (2.7) and, therefore, it also implies condition (2.18).
Note 2.5.
In the SRS-based context, that is, when $F_{n}$ is $F_{n,\textsc{srs}}$, Shao (2003) derives exponential bounds for the probability $\mathbb{P}(|F_{n,\textsc{srs}}^{-1}(p)-F^{-1}(p)|\geq\varepsilon)$ with the aim at establishing strong consistency. We are concerned with (weak) consistency and so do not need exponential bounds. In other words, we just need convergence of $\mathbb{P}(|F_{n}^{-1}(p)-F^{-1}(p)|\geq\varepsilon)$ to 00 for every $\varepsilon>0$ when $n\to\infty$, and for this, condition (2.19) imposed on the generic sequence $F_{n}$, $n\in\mathbb{N}$, is sufficient.
In Section 3 we shall give an illustrative numerical study showing that statement (2.18) fails for a cdf $F$ for which the quantile function $F^{-1}$ is discontinuous at a point $p$. The same example will also show that the empirical upper-layer integral $\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ fails asymptotic normality. This phenomenon is natural given the results of Stigler (1973); see also Gribkova and Helmers (2011) for further notes on the topic.
Example 2.3 (SRS).
When $F_{n}$ is $F_{n,\textsc{srs}}$, condition (2.19) is a consequence of the Kolmogorov-Smirnov theorem. Since condition (2.19) is satisfied, condition (2.18) is satisfied as well, provided that, according to Lemma 2.1, the quantile function $F^{-1}$ is continuous at the point $p$. Furthermore, since the integral $\int^{\infty}_{F^{-1}(p)}\big{(}1-F_{n,\textsc{srs}}(x)\big{)}\mathrm{d}x$ is the arithmetic mean of $n$ independent copies of random variable (2.13), we conclude that the integral $\int^{\infty}_{F^{-1}(p)}\big{(}1-F_{n,\textsc{srs}}(x)\big{)}\mathrm{d}x$ satisfies the central limit theorem provided that random variable (2.13) has a finite second moment. This is so provided that $F\in\mathcal{L}^{+}_{2}$ (see Lemma A.1 for details), where $\mathcal{L}^{+}_{2}$ denotes the set of all cdf’s such that the integral $\int_{p}^{1}\big{(}F^{-1}(u)\big{)}^{2}\mathrm{d}u$ is finite. Note that $F\in\mathcal{L}^{+}_{2}$ is equivalent to saying that the positive part $X^{+}$ of the random variable $X\sim F$ has a finite second moment $\mathbb{E}((X^{+})^{2})<\infty$. In summary, when $F\in\mathcal{L}^{+}_{2}$ and the quantile function $F^{-1}$ is continuous at the point $p$, we have the asymptotic normality result
$\sqrt{n}\left(\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u-\int_{p}^{1}F^{-1}(u)\mathrm{d}u\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}(0,\sigma_{\textsc{upper}}^{2}(p)),$ (2.23)
where $\sigma_{\textsc{upper}}^{2}(p)$ is the variance of the random variable $Y_{\textsc{upper}}(p)$. Note that the variance can be expressed in terms of the cdf $F$ by the formula
$\sigma_{\textsc{upper}}^{2}(p)=\int^{\infty}_{F^{-1}(p)}\int^{\infty}_{F^{-1}(p)}\big{(}F(x\wedge y)-F(x)F(y)\big{)}\mathrm{d}x\mathrm{d}y,$ (2.24)
although we have found that the following alternative representations (recall equation (2.13))
$\sigma_{\textsc{upper}}^{2}(p)$ $=\mathrm{Var}\left(\big{(}X-F^{-1}(p)\big{)}^{+}\right)$
$=\mathrm{Var}\left(\big{(}F^{-1}(U)-F^{-1}(p)\big{)}^{+}\right)$ (2.25)
are sometimes easier to use, where $U$ denotes a uniform on $[0,1]$ random variable.
Note 2.6 (existence of uniform rv’s).
[…]
Note 2.7 (time series).
Quite often in applications, data arrive in the form of time series, and thus researchers have been interested in deriving large sample properties for various tail risk measures based on such data (e.g., Davydov and Zitikis, 2002, 2003; Helmers and Zitikis, 2005; Davydov et al., 2007; Chen, 2008; Linton and Xiao, 2013; Ziegel et al., 2020; Moutanabbir and Bouaddi, 2024; García-Risueño, 2025, and references therein). Corollaries 2.1 and 2.3, which are not attached to any particular sampling design or dependence structure, are applicable in such scenarios. Indeed, statements like those in condition (2.7) for consistency and condition (2.19) for asymptotic distribution have been verified for various classes of time series, where the empirical cdf plays the role of $F_{n}$. To give some flavour of what is known, we first note that in the case of weakly-dependent stationary time series, the normalizing constant $A_{n}$ is $\sqrt{n}$ and the limiting process of $A_{n}(F_{n}-F)$ is normal, whereas in the case of long-memory stationary time series, the normalizing constant $A_{n}$ is of the order $n^{\theta}$ for various $\theta\in(0,1/2)$ and the limiting process, if exists, is non-normal. The monographs by Dehling et al. (2002), Rio (2017), and Doukhan (2018) offer a wealth of information on the topic.
We next present convenient computational formulas for the SRS-based empirical upper-layer integral $\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$.
Computation 2.1 (SRS).
When computing, estimators in the form of integrals are not convenient to work with. Hence, using equation (2.12), we shall now express $\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ in terms of the order statistics $X_{1:n}\leq\cdots\leq X_{n:n}$, whose realized values are known in practice. Namely, we have
$\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ $=\int_{p}^{\lceil np\rceil/n}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u+\sum_{i=\lceil np\rceil+1}^{n}\int_{(i-1)/n}^{i/n}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$
[…]
Equations (2.26) and (2.27) give computationally-friendly formulas for the empirical upper-layer integral, which is connected to the left-trimmed mean (e.g., Stigler, 1973) via equation (2.27).
3 The upper-layer integral: simulated experiments
In this section we present several experiments that empirically illustrate some of the theoretical considerations in previous Section 2.
Simulation 3.1 (SRS; bias).
We shall now empirically illustrate that $\mathrm{Bias}_{n,\textsc{srs}}^{\textsc{upper}}(p)$ is non-positive, as stated in Example 2.2. To avoid confounding negativity of bias with skewness of distribution, we work with a symmetric (around its mean) and light-tailed distribution. Namely, let $V$ be the uniform on $[0,2]$ random variable, whose quantile function is $F^{-1}(u)=2u$ for $0\leq u\leq 1$. We fix the lower-layer probability level
$p={3\over 4}$
and choose several values of $n$ (to be specified in a moment) to get an insight into the behaviour of the bias with respect to the sample size. The population upper-layer integral $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$ is equal to $161-p^{2}=7/$ 16. We obtain the first SRS by observing $n$ independent copies $V_{1},\dots,V_{n}$ of $V$, and then calculate the empirical upper-layer integral $\int_{p}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ according to formula (2.26). This gives us the first difference
[…]
$n$ 40 100 200 500 1000
$B^{\textsc{ave}}_{n}$0.004855 -0.002113 -0.000944 -0.000311 -0.000154
$B^{\textsc{med}}_{n}$0.002482 -0.001203 -0.000463 -0.000207 -0.000099
[…]
we report the values of $B^{\textsc{ave}}_{n}$ and $B^{\textsc{med}}_{n}$ for different $n$’s. Since $m$ is very large, we expect the averages $B^{\textsc{ave}}_{n}$ reported in Table 3.1 to be close to the corresponding theoretical biases $\mathrm{Bias}_{n,\textsc{srs}}^{\textsc{upper}}$, which are defined by equation (2.17). Indeed, the table shows that all $B^{\textsc{ave}}_{n}$’s are negative. Furthermore, their absolute values are decreasing when $n$ grows, which is expected given the proven consistency of the empirical upper-layer integral. Nevertheless, as a follow-up research topic, it is worth thinking of modifying the empirical upper-layer integral to reduce its bias. Finally, we note that the reason behind presenting the medians $B^{\textsc{med}}_{n}$ in Table 3.1 is that by comparing them with the averages $B^{\textsc{ave}}_{n}$, we can get insights about the skewness of the distribution of individual biases.
Simulation 3.2 (SRS; asymptotic normality holds).
Let $V$ be a uniform on $[0,2]$ random variable, and so its cdf is $F(x)=x/2$ for all $0\leq x\leq 2$. We set the parameter $p$ value to
$p={1\over 2}$
[…]
$\Delta_{n}:={\sqrt{n}\over\sigma_{\textsc{upper}}(1/2)}\left(\int_{1/2}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u-\int_{1/2}^{1}F^{-1}(u)\mathrm{d}u\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1),$ (3.1)
where
$\int_{1/2}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ $={1\over n}\sum_{i=(n/2)+1}^{n}V_{i:n}\quad\big{(}\textrm{because$\lceil n/2\rceil=n/2$}\big{)},$
$\int_{1/2}^{1}F^{-1}(u)\mathrm{d}u$ $={3\over 4},$
$\sigma_{\textsc{upper}}^{2}(1/2)$ $={5\over 48}.$ (3.2)
[…]
which supports the limiting standard normal distribution claimed in statement (3.1).
Simulation 3.3 (SRS; asymptotic normality fails).
To illustrate the pivotal role of continuity of the quantile function $F^{-1}$ at the point $p$ and thus of condition (2.18) in the statement of asymptotic normality of the empirical upper-layer integral, we next introduce a new random variable, denoted by $Z$. Its definition is based on a uniform on $[0,2]$ random variable $V$. Namely, we fix any constant $a\in[0,1]$ and define a “gapped” random variable $Z$ by the formula
$Z=(1-a)V\mathds{1}\{V\leq 1\}+\big{(}2a+(1-a)V\big{)}\mathds{1}\{V>1\}.$
Note that when $a=0$, the gap vanishes and $Z$ turns into the original uniform random variable $V$, the case that we have discussed in previous Simulation 3.2. When $a=1$, the random variable $Z$ turns into the discrete random variable taking values 00 and $2$ with probabilities $1/2$. To better understand and appreciate the random variable $Z$ with respect to the parameter $a\in(0,1)$, we note the following properties of $Z$ when it is viewed as a function of the argument $V$:
  • if $V=0$, then $Z=0$;
  • if $V=1$, then $Z=1-a$;
  • if $V\downarrow 1$, then $Z\downarrow 1+a$;
  • if $V=2$, then $Z=2$;
Hence, all the values of $Z$ are in the interval $[0,2]$, although there are no values in $(1-a,1+a]$, which represents the gap in the distribution of $Z$. The cdf $z\mapsto H(z)$ of $Z$, which is visualized in Figure 3.2,
(a) The cdf $H$.
(b) The quantile function $H^{-1}$.
Figure 3.2: The distribution of the gapped random variable $Z$.
is strictly increasing at every point $z\in(0,2)\setminus[1-a,1+a]$ and is constant on the interval $[1-a,1+a]$. The quantile $H^{-1}(1/2)$, which is the median of $Z$, is the left-hand endpoint $1-a$ of the gap of $Z$. Consequently, the quantile function $H^{-1}$ is not continuous at the point $p=1/2$. This helps us to understand the reason why in the case
$p={1\over 2},$ (3.3)
we have the following two non-convergence statements: first, non-consistency
$H_{n,\textsc{srs}}^{-1}(1/2)\stackrel{{\scriptstyle\mathbb{P}}}{{\nrightarrow}}H^{-1}(1/2),$ (3.4)
and, second, asymptotic non-normality
$\Delta_{n}:={\sqrt{n}\over\sigma_{\textsc{upper}}(1/2)}\left(\int_{1/2}^{1}H_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u-\int_{1/2}^{1}H^{-1}(u)\mathrm{d}u\right)\stackrel{{\scriptstyle d}}{{\nrightarrow}}\mathcal{N}(0,1),$ (3.5)
where (recall equation (2.24))
$\sigma_{\textsc{upper}}^{2}(1/2)=\int^{2}_{1-a}\int^{2}_{1-a}\big{(}H(x\wedge y)-H(x)H(y)\big{)}\mathrm{d}x\mathrm{d}y.$ (3.6)
Although statements (3.4) and (3.5) can be verified theoretically, which we shall discuss in a moment, we find it illuminating to demonstrate the two statements empirically. For this, we set the gap radius to
$a={1\over 2}.$
Hence, the distribution of $Z$ has the gap $[1/2,3/2]$ in its support $[0,2]$. When simulating we conveniently work with even sample sizes $n$ and thus have the equation $\lceil n/2\rceil=n/2$, which simplifies formulas of the empirical estimators, especially of the upper-layer integral. We proceed by noting that given the nature of the random variable $V$, which is symmetric around its mean $1$, and given the definition of the gapped $Z$, we expect to see approximately a half of the simulated $Z$ values below $1/2$ and the remaining ones above $3/2$. Indeed, when $n\to\infty$, we have the following two statements (see Appendix A.6.2)
$\mathbb{P}\left(Z_{\lceil n/2\rceil:n}>3/2\right)\to 1/2,$ (3.7)
$\mathbb{P}\left(Z_{\lceil n/2\rceil:n}<1/2\right)\to 1/2.$ (3.8)
This explains non-consistency statement (3.4), which we have visualized in Figure 3.3
Figure 3.3: Simulated $m=100$ values of $H_{n,\textsc{srs}}^{-1}(1/2)$ for various sample sizes $n$.
using the formulas
$H_{n,\textsc{srs}}^{-1}(1/2)$ $=Z_{\lceil n/2\rceil:n}\leavevmode\nobreak\ \leavevmode\nobreak\ (=Z_{(n/2):n}),$
$H^{-1}(1/2)$ $=1/2$
for various (even) sample sizes $n$. As to asymptotic non-normality statement (3.5), we have used the formulas
$\int_{1/2}^{1}H_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ $={1\over n}\sum_{i=(n/2)+1}^{n}Z_{i:n}\quad\big{(}\textrm{because$\lceil n/2\rceil=n/2$}\big{)},$
$\int_{1/2}^{1}H^{-1}(u)\mathrm{d}u$ $={7\over 8},$
$\sigma_{\textsc{upper}}^{2}(1/2)$ $={77\over 192}.$ (3.9)
Figure 3.4
[…]
Note 3.1.
When calculating the variance $\sigma_{\textsc{upper}}^{2}(1/2)$ in equation (3.9), we found that the quantile-based expression (recall equation (2.25))
$\sigma_{\textsc{upper}}^{2}(1/2)$ $=\mathrm{Var}\left(\big{(}H^{-1}(U)-H^{-1}(1/2)\big{)}^{+}\right)$
$=\mathrm{Var}\left(\big{(}H^{-1}(U)-(1-a)\big{)}^{+}\right)$ (3.10)
is a particularly convenient starting point, where $U$ denotes a uniform on $[0,1]$ random variable, with $1/2$ referring to our choice of $p$ made in (3.3). In this way, we also easily arrive at the expression
${29a^{2}+14a+5\over 48}$ (3.11)
for $\sigma_{\textsc{upper}}^{2}(1/2)$ in the case of arbitrary $a\in[0,1]$ (see Appendix A.6.3 for details). Hence, $\sigma_{\textsc{upper}}^{2}(1/2)=5/48$ when $a=0$ (equation (3.2)), $\sigma_{\textsc{upper}}^{2}(1/2)=77/192$ when $a=1/2$ (equation (3.9)), and $\sigma_{\textsc{upper}}^{2}(1/2)=1$ when $a=1$ (in which case $Z$ turns into a discrete random variable taking values 00 and $2$ with probabilities $1/2$). Of course, we could have also calculated $\sigma_{\textsc{upper}}^{2}(1/2)$ using equation (3.6) but this cdf-based route is more involved.
4 The lower-layer integral
In this section we develop statistical inference for the lower-layer integral $\int_{0}^{p}F^{-1}(u)\mathrm{d}u$ for any fixed upper probability level
$p\in(0,1).$
Naturally, we assume that the cdf $F$ is from the class $\mathcal{L}^{-}_{1}$ of those cdf’s for which the integral $\int_{0}^{p}F^{-1}(u)\mathrm{d}u$ is finite, which is the same as to require that the negative part $X^{-}=(-X)\vee 0$ of the random variable $X\sim F$ has a finite first moment $\mathbb{E}(X^{-})<\infty$.
Note 4.1.
When progressing through this section, we shall see that the narrative closely follows that of Section 2. This may prompt us to think that the results could be derived from those of Section 2 by simply replacing $X$ by $-X$. This train of thought, though useful, is fraught with danger because $F^{-1}$ is not the ordinary inverse of $F$ and so subtleties do arise. It is our desire, therefore, to present the results in such a way that the user would be able to use them directly, without needing to work out the underlying theory.
We continue our analysis of the lower-layer integral by noting that the functional
$\mathcal{L}^{-}_{1}\ni F\mapsto\int_{0}^{p}F^{-1}(u)\mathrm{d}u\in\mathbb{R}$
is not linear. It appears that when developing statistical inference it can be approximated by the linear one
$\mathcal{L}^{-}_{1}\ni G\mapsto\int_{-\infty}^{F^{-1}(p)}G(x)\mathrm{d}x\in\mathbb{R}.$ (4.1)
Indeed, we shall see from the following corollaries that the difference between the empirical lower-layer integral $\int_{0}^{p}F_{n}^{-1}(u)\mathrm{d}u$ and its population counterpart $\int_{0}^{p}F^{-1}(u)\mathrm{d}u$ gets asymptotically close to the difference between the integrals $\int_{-\infty}^{F^{-1}(p)}F(x)\mathrm{d}x$ and $\int_{-\infty}^{F^{-1}(p)}F_{n}(x)\mathrm{d}x$, where $F_{n}\in\mathcal{L}^{-}_{1}$, $n\in\mathbb{N}$, are cdf’s approaching $F$ when the parameter $n$ grows indefinitely. The corollaries rely on the following theorem, whose proof is given in Appendix A.2.
Theorem 4.1.
Let $F$ and $G$ be any two cdf’s from $\mathcal{L}^{-}_{1}$. Then
$\int_{0}^{p}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u=\int_{-\infty}^{F^{-1}(p)}\big{(}F(x)-G(x)\big{)}\mathrm{d}x+\mathrm{Rem}(p;F,G),$ (4.2)
where the non-negative remainder term $\mathrm{Rem}(p;F,G)$ is defined by equation (2.3) and satisfies bounds (2.4) and (2.5).
We are now ready to formulate and discuss several corollaries to Theorem 4.1.
Corollary 4.1 (consistency).
Suppose that $F\in\mathcal{L}^{-}_{1}$, and let $F_{1},F_{2},\ldots\in\mathcal{L}^{-}_{1}$ be any sequence of cdf’s satisfying condition (2.7). Then
$\int_{0}^{p}F_{n}^{-1}(u)\mathrm{d}u-\int_{0}^{p}F^{-1}(u)\mathrm{d}u=\int_{-\infty}^{F^{-1}(p)}\big{(}F(x)-F_{n}(x)\big{)}\mathrm{d}x+o_{\mathbb{P}}(1)$ (4.3)
and, therefore, the consistency statement
$\int_{0}^{p}F_{n}^{-1}(u)\mathrm{d}u\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int_{0}^{p}F^{-1}(u)\mathrm{d}u$ (4.4)
holds if and only if
$\int_{-\infty}^{F^{-1}(p)}F_{n}(x)\mathrm{d}x\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int_{-\infty}^{F^{-1}(p)}F(x)\mathrm{d}x.$ (4.5)
Proof.
The corollary follows from Theorem 4.1 and bound (2.5) with $F_{n}$ instead of $G$, because $F^{-1}(p)$ is finite and $F_{n}^{-1}(p)$ is asymptotically bounded, and therefore condition (2.7) implies $\mathrm{Rem}(p;F,G)=o_{\mathbb{P}}(1)$ due to bound (2.5). This establishes statement (4.3), which implies the equivalence of statements (4.4) and (4.5), and concludes the proof of Corollary 4.1. ∎
Example 4.1 (SRS).
Condition (2.7) with $F_{n,\textsc{srs}}$ instead of $F_{n}$ is satisfied by the Glivenko-Cantelli theorem. Furthermore, by linearity of functional (4.1), the integral $\int_{-\infty}^{F^{-1}(p)}F_{n}(x)\mathrm{d}x$ is the arithmetic mean $n^{-1}\sum_{i=1}^{n}Y_{i,\textsc{lower}}$ of $n$ independent copies of the random variable
$Y_{\textsc{lower}}(p)$ $=\int_{-\infty}^{F^{-1}(p)}\mathds{1}\{X\leq x\}\mathrm{d}x$
$=\big{(}F^{-1}(p)-X\big{)}^{+}.$ (4.6)
The random variable $Y_{\textsc{lower}}(p)$ has a finite first moment because $F\in\mathcal{L}^{-}_{1}$ (see Lemma A.2 for details), and so statement (4.5) with $F_{n,\textsc{srs}}$ instead of $F_{n}$ holds. By Corollary 4.1, therefore, the empirical lower-level integral $\int_{0}^{p}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ is a consistent estimator of $\int_{0}^{p}F^{-1}(u)\mathrm{d}u$. In summary, when $F\in\mathcal{L}^{-}_{1}$, we have
$\int_{0}^{p}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int_{0}^{p}F^{-1}(u)\mathrm{d}u.$ (4.7)
Note 4.2 (dependent data).
Although we work with the case $p\in(0,1)$, let us set $p=1$ for a moment. In this case, statement (4.7) turns into the law of large numbers
$\mu_{n,\textsc{srs}}=\int_{0}^{1}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\mu,$ (4.8)
which of course holds under SRS. A departure from this sampling design leads to ergodic theory, and in particular to sequences $X_{1},X_{2},\dots$ that are stationary and ergodic. Hence, it is only natural that results of the type of statement (4.7) when $p\in(0,1)$ have been established within the class of stationary and ergodic sequences (e.g., Aaronson et al., 1996; Gilat and Helmers, 1997; Davydov and Zitikis, 2002; Helmers and Zitikis, 2005, and references therein). It is also useful to point out that while the latter two references deal with the integral $\int_{0}^{p}F^{-1}(u)\mathrm{d}u$, the first two references deal with the more general integral $\int_{0}^{1}F^{-1}(u)w(u)\mathrm{d}u$, which we shall discuss in more detail in Section 7.
Corollary 4.2 (bias).
Suppose that $F\in\mathcal{L}^{-}_{1}$, and let $F_{1},F_{2},\ldots\in\mathcal{L}^{-}_{1}$ be any sequence of cdf’s that are unbiased estimators of $F$, that is, satisfy condition (2.15), and such that condition (2.16) is satisfied. Then $\int_{0}^{p}F_{n}^{-1}(u)\mathrm{d}u$ is a non-negatively biased estimator of $\int_{0}^{p}F^{-1}(u)\mathrm{d}u$.
Proof.
Using Theorem 4.1 with $F_{n}$ instead of $G$, and also recalling that the remainder term $\mathrm{Rem}(p;F,G)$ is non-negative, we have
$\mathrm{Bias}_{n}^{\textsc{lower}}(p):=$ $\mathbb{E}\bigg{(}\int_{0}^{p}F_{n}^{-1}(u)\mathrm{d}u\bigg{)}-\int_{0}^{p}F^{-1}(u)\mathrm{d}u$
$=$ $\mathbb{E}\big{(}\mathrm{Rem}(p;F,F_{n})\big{)}\in[0,\infty)$
for every $n\in\mathbb{N}$. Note that $\mathbb{E}\big{(}\mathrm{Rem}(p;F,F_{n})\big{)}$ is finite because of bound (2.5) and condition (2.16). This establishes Corollary 4.2. ∎
Example 4.2 (SRS).
When $F_{n}$ is $F_{n,\textsc{srs}}$, condition (2.15) is satisfied. Furthermore, condition (2.16) is satisfied when $F\in\mathcal{T}_{\varepsilon}$ for some $\varepsilon>0$ (recall Example 2.2 for details). Hence, we conclude from Corollary 4.2 that when $F\in\mathcal{L}^{-}_{1}\cap\mathcal{T}_{\varepsilon}$, the empirical estimator $\int_{0}^{p}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ of the lower-layer integral $\int_{0}^{p}F^{-1}(u)\mathrm{d}u$, although being consistent by Example 4.1, has the positive bias
$\mathrm{Bias}_{n,\textsc{srs}}^{\textsc{lower}}(p):=\mathbb{E}\big{(}\mathrm{Rem}(p;F,F_{n,\textsc{srs}})\big{)}\in[0,\infty).$ (4.9)
Note 4.3.
The condition that $F\in\mathcal{L}^{-}_{1}\cap\mathcal{T}_{\varepsilon}$ for some $\varepsilon>0$ can equivalently be reformulated as $F\in\mathcal{L}^{-}_{1}\cap\mathcal{T}^{+}_{\varepsilon}$ for some $\varepsilon>0$, where $\mathcal{T}^{+}_{\varepsilon}$ consists of all those cdf’s that satisfy $x^{\varepsilon}(1-F(x))\to 0$ when $x\to\infty$. This reformulation is possible due to the fact that $F\in\mathcal{L}^{-}_{1}$ implies $|x|^{\varepsilon}F(x)\to 0$ when $x\to-\infty$ for every $\varepsilon\in(0,1)$, and we need this condition to hold for just one $\varepsilon>0$. For those researchers who think in terms of moments, a sufficient condition that ensures the existence of $\varepsilon>0$ such that $F\in\mathcal{L}^{-}_{1}\cap\mathcal{T}^{+}_{\varepsilon}$ would be the requirement that
$\mathbb{E}(X^{-})<\infty\quad\text{and}\quad\mathbb{E}((X^{+})^{\varepsilon})<\infty$
for some $\varepsilon>0$, where $X\sim F$. Note that this requirement can equivalently be rewritten as $F\in\mathcal{L}^{-}_{1}\cap\mathcal{L}^{+}_{\varepsilon}$.
Note 4.4.
Let $F\in\mathcal{L}_{1}$, where $\mathcal{L}_{1}:=\mathcal{L}^{-}_{1}\cap\mathcal{L}^{+}_{1}$. Hence, $F\in\mathcal{L}_{1}$ is equivalent to saying that $\int_{0}^{1}F^{-1}(u)\mathrm{d}u\in\mathbb{R}$, which is the same as saying that the random variable $X\sim F$ has a finite first moment $\mathbb{E}(X)\in\mathbb{R}$. When in addition to this moment requirement, condition (2.15) is also satisfied, we have the equation
$\mathrm{Bias}_{n}^{\textsc{lower}}(p)=-\mathrm{Bias}_{n}^{\textsc{upper}}(p)$
because, due to equation (1.1) with $F_{n}$ instead of $G$,
$\mathrm{Bias}_{n}^{\textsc{upper}}(p)+\mathrm{Bias}_{n}^{\textsc{lower}}(p)$ $=\mathbb{E}\bigg{(}\int_{0}^{1}F_{n}^{-1}(u)\mathrm{d}u-\int_{0}^{1}F^{-1}(u)\mathrm{d}u\bigg{)}$
$=\mathbb{E}\bigg{(}\int^{\infty}_{-\infty}\big{(}F(x)-F_{n}(x)\big{)}\mathrm{d}x\bigg{)}$
$=\int^{\infty}_{-\infty}\mathbb{E}\big{(}F(x)-F_{n}(x)\big{)}\mathrm{d}x,$
where the right-hand side is equal to 00 due to condition (2.15).
Corollary 4.3.
Suppose that $F\in\mathcal{L}^{-}_{1}$, and let $F_{1},F_{2},\ldots\in\mathcal{L}^{-}_{1}$ be any sequence of cdf’s satisfying condition (2.18) and also condition (2.19) with some normalizing constants $A_{n}\to\infty$ when $n\to\infty$. Then
$A_{n}\left(\int_{0}^{p}F_{n}^{-1}(u)\mathrm{d}u-\int_{0}^{p}F^{-1}(u)\mathrm{d}u\right)=A_{n}\int_{-\infty}^{F^{-1}(p)}\big{(}F(x)-F_{n}(x)\big{)}\mathrm{d}x+o_{\mathbb{P}}(1)$ (4.10)
and, therefore, the convergence-in-distribution statement
$A_{n}\left(\int_{0}^{p}F_{n}^{-1}(u)\mathrm{d}u-\int_{0}^{p}F^{-1}(u)\mathrm{d}u\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}_{\textsc{lower}}(p)$ (4.11)
holds if and only if
$A_{n}\left(\int_{-\infty}^{F^{-1}(p)}F_{n}(x)\mathrm{d}x-\int_{-\infty}^{F^{-1}(p)}F(x)\mathrm{d}x\right)\stackrel{{\scriptstyle d}}{{\to}}-\mathcal{L}_{\textsc{lower}}(p),$ (4.12)
where $\mathcal{L}_{\textsc{lower}}(p)$ is a random variable determined by statement (4.12).
Proof.
Using Theorem 4.1 with $F_{n}$ instead of $G$, we only need to show that $A_{n}\mathrm{Rem}(p;F,F_{n})$ converges in probability to 00, which we showed in the proof of Corollary 2.3. This establishes equation (4.10), which implies the equivalence of statements (4.11) and (4.12), and concludes the proof of Corollary 4.3. ∎
Example 4.3 (SRS).
When $F_{n}$ is $F_{n,\textsc{srs}}$, condition (2.19), which is a requirement in Corollary 4.3, is satisfied by the Kolmogorov-Smirnov theorem. Hence, condition (2.18) is satisfied as well, provided that, according to Lemma 2.1, the quantile function $F^{-1}$ is continuous at the point $p$. Furthermore, since the integral $\int_{-\infty}^{F^{-1}(p)}F_{n}(x)\mathrm{d}x$ is the arithmetic mean of $n$ independent copies of random variable (4.1), we conclude that $\int_{-\infty}^{F^{-1}(p)}F_{n}(x)\mathrm{d}x$ satisfies the central limit theorem provided that random variable (4.1) has a finite second moment. This is so provided that $F\in\mathcal{L}^{-}_{2}$ (see Lemma A.2 for details), where $\mathcal{L}^{-}_{2}$ denotes the set of all cdf’s such that the integral $\int_{0}^{p}\big{(}F^{-1}(u)\big{)}^{2}\mathrm{d}u$ is finite. Note that $F\in\mathcal{L}^{-}_{2}$ is equivalent to saying that the negative part $X^{-}$ of the random variable $X\sim F$ has a finite second moment $\mathbb{E}((X^{-})^{2})<\infty$. In summary, when $F\in\mathcal{L}^{-}_{2}$ and the quantile function $F^{-1}$ is continuous at the point $p$, we have the asymptotic normality result
$\sqrt{n}\left(\int_{0}^{p}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u-\int_{0}^{p}F^{-1}(u)\mathrm{d}u\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}(0,\sigma_{\textsc{lower}}^{2}(p)),$ (4.13)
where $\sigma_{\textsc{lower}}^{2}(p)$ is the variance of the random variable $Y_{\textsc{lower}}(p)$. The variance can be expressed by the formula
$\sigma_{\textsc{lower}}^{2}(p)=\int_{-\infty}^{F^{-1}(p)}\int_{-\infty}^{F^{-1}(p)}\big{(}F(x\wedge y)-F(x)F(y)\big{)}\mathrm{d}x\mathrm{d}y.$ (4.14)
Note 4.5 (dependent data).
Limiting distributions extending statement (4.13) to classes of dependent random sequences were studied by Davydov and Zitikis (2003), and Davydov et al. (2007). The theory of empirical processes based on such random sequences becomes particulary useful, and we refer to Dehling et al. (2002) for further details on the topic. Without going into the complexities of such studies, we note that dependence structures modify the asymptotic variance, and even the asymptotic distribution, which may become more complex than the mere normal distribution.
Computation 4.1 (SRS).
Using equation (2.12) for calculating the quantile $F_{n,\textsc{srs}}^{-1}(u)$, we have
$\int_{0}^{p}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ $=\sum_{i=1}^{\lceil np\rceil}\int_{(i-1)/n}^{i/n}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u-\int_{p}^{\lceil np\rceil/n}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$
[…]
These expressions give computationally-friendly formulas for the empirical lower-layer integral, which is connected to the right-trimmed mean (e.g., Stigler, 1973) via equation (4.15).
5 The middle-layer integral
Let $\mathcal{L}$ denote the class of all cdf’s. Fix any pair of probability levels $p_{1}$ and $p_{2}$ such that
$10<p_{1}<p_{2}<$ 1 .
The middle-layer integral $\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u$ is finite for every $F\in\mathcal{L}$. The functional
$\mathcal{L}\ni F\mapsto\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u\in\mathbb{R}$
is not linear, but when developing statistical inference it can be approximated by the linear one
$\mathcal{L}\ni G\mapsto\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}G(x)\mathrm{d}x\in\mathbb{R}.$ (5.1)
Indeed, we shall see from the following corollaries that the difference between the empirical middle-layer integral $\int_{p_{1}}^{p_{2}}F_{n}^{-1}(u)\mathrm{d}u$ and its population counterpart $\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u$ gets asymptotically close to the difference between the integrals $\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F(x)\mathrm{d}x$ and $\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F_{n}(x)\mathrm{d}x$, where $F_{n}$, $n\in\mathbb{N}$, are cdf’s approaching $F$ when the parameter $n$ grows indefinitely. The corollaries rely on the following theorem, whose proof is given in Appendix A.2.
Theorem 5.1.
Let $F$ and $G$ be any two cdf’s. Then
$\int_{p_{1}}^{p_{2}}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u=\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}F(x)-G(x)\big{)}\mathrm{d}x+\mathrm{Rem}(p_{2};F,G)-\mathrm{Rem}(p_{1};F,G),$ (5.2)
where the non-negative remainder terms $\mathrm{Rem}(p_{1};F,G)$ and $\mathrm{Rem}(p_{2};F,G)$ are defined by equation (2.3) and satisfy bounds (2.4) and (2.5).
Note 5.1.
It is tempting to collapse equations (1.1), (2.2), (4.2), and (5.2) into one by relaxing the restriction $10<p_{1}<p_{2}<$ 1 to $0\leq p_{1}<p_{2}\leq 1$. This is indeed possible by augmenting the definition of the remainder term $\mathrm{Rem}(p;F,G)$, which has so far been given only for $p\in(0,1)$, by setting $\mathrm{Rem}(p;F,G)$ to 00 when $p\in\{0,1\}$. If we agree with this augmentation, then it is also imperative to replace $F^{-1}(p_{1})$ by $\infty$- ∞ when $p_{1}=0$ in the integral
$\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}F(x)-G(x)\big{)}\mathrm{d}x$ (5.3)
on the right-hand side of equation (5.2), and to also replace $F^{-1}(p_{2})$ by $\infty$ when $p_{2}=1$. Note that $F^{-1}(0)$ may or may not be equal to $\infty$- ∞, and $F^{-1}(1)$ may or may not be equal to $\infty$. Hence, although the noted replacements may look artificial, they are absolutely necessary, as next Example 5.1 shows. Consequently, the role of infinities $\infty$- ∞ and $\infty$ instead of $F^{-1}(p_{1})$ and $F^{-1}(p_{2})$ in integral (5.3) when $p_{1}=0$ and $p_{2}=1$, respectively, is crucial. In summary, we recommend having four separate equations (1.1), (2.2), (4.2), and (5.2) in our toolbox, instead of having just one general equation with a number of caveats and adjustments. The four separate equations should help to avoid potential overlooks and pitfalls.
Example 5.1.
Let $F\sim U_{[0,1]}$ and $G\sim U_{[0,2]}$ be two random variables with uniform on the intervals $[0,1]$ and $[0,2]$ distributions, respectively. That is, $F(x)=x$ when $0\leq x\leq 1$ and $G(x)=x/2$ when $0\leq x\leq 2$. Consequently, $F^{-1}(t)=t$ and $G^{-1}(t)=2t$ when $0\leq t\leq 1$. From these formulas, we obtain
$\int_{0}^{1}\big{(}G^{-1}(t)-F^{-1}(t)\big{)}\mathrm{d}t={1\over 2},$
$\int_{F^{-1}(0)}^{F^{-1}(1)}\big{(}F(x)-G(x)\big{)}\mathrm{d}x={1\over 4},$
$\int_{-\infty}^{\infty}\big{(}F(x)-G(x)\big{)}\mathrm{d}x={1\over 2}.$
Hence,
$\int_{0}^{1}\big{(}G^{-1}(t)-F^{-1}(t)\big{)}\mathrm{d}t=\int_{-\infty}^{\infty}\big{(}F(x)-G(x)\big{)}\mathrm{d}x,$
which is what equation (1.1) says, but
$\int_{0}^{1}\big{(}G^{-1}(t)-F^{-1}(t)\big{)}\mathrm{d}t\neq\int_{F^{-1}(0)}^{F^{-1}(1)}\big{(}F(x)-G(x)\big{)}\mathrm{d}x.$
This concludes Example 5.1.
We are now ready to formulate and discuss several corollaries to Theorem 5.1.
Corollary 5.1 (consistency).
Let $F$ be any cdf, and let $F_{1},F_{2},\ldots$ be any sequence of cdf’s satisfying condition (2.7). Then
$\int_{p_{1}}^{p_{2}}F_{n}^{-1}(u)\mathrm{d}u-\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u=\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}F(x)-F_{n}(x)\big{)}\mathrm{d}x+o_{\mathbb{P}}(1)$ (5.4)
and, therefore, the consistency statement
$\int_{p_{1}}^{p_{2}}F_{n}^{-1}(u)\mathrm{d}u\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u$ (5.5)
holds if and only if
$\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F_{n}(x)\mathrm{d}x\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F(x)\mathrm{d}x.$ (5.6)
Proof.
The corollary follows from Theorem 5.1 and bound (2.5) with $F_{n}$ instead of $G$, because for $p\in\{p_{1},p_{2}\}\subset(0,1)$, $F^{-1}(p)$ is finite and $F_{n}^{-1}(p)$ is asymptotically bounded. Therefore condition (2.7) implies $\mathrm{Rem}(p;F,F_{n})=o_{\mathbb{P}}(1)$ for both $p=p_{1}$ and $p=p_{2}$, and thus, in turn, implies statement (5.4). This establishes the equivalence of statements (5.5) and (5.6), and concludes the proof of Corollary 5.1. ∎
Example 5.2 (SRS).
Condition (2.7) with $F_{n,\textsc{srs}}$ in place of $F_{n}$ is satisfied by the Glivenko-Cantelli theorem. Furthermore, by linearity of functional (5.1), the integral $\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F_{n}(x)\mathrm{d}x$ is the arithmetic mean $n^{-1}\sum_{i=1}^{n}Y_{i,\textsc{middle}}(p_{1},p_{2})$ of $n$ independent copies of the random variable
$Y_{\textsc{middle}}(p_{1},p_{2})$ $=\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\mathds{1}\{X\leq x\}\mathrm{d}x$
$=\big{(}F^{-1}(p_{2})-X\big{)}^{+}-\big{(}F^{-1}(p_{1})-X\big{)}^{+}.$ (5.7)
The random variable $Y_{\textsc{middle}}(p_{1},p_{2})$ always has a finite first moment (see Note A.1 for details). Consequently, $\int_{p_{1}}^{p_{2}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ is a consistent estimator of $\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u$, that is, for every cdf $F$, we have
$\int_{p_{1}}^{p_{2}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u.$ (5.8)
Corollary 5.2 (bias).
Let $F$ be any cdf, and let $F_{1},F_{2},\ldots$ be any sequence of cdf’s that are unbiased estimators of $F$, that is, satisfy condition (2.15), and such that condition (2.16) is satisfied with $p\in\{p_{1},p_{2}\}$. Then
$\mathrm{Bias}_{n}^{\textsc{middle}}(p_{1},p_{2}):=$ $\mathbb{E}\bigg{(}\int_{p_{1}}^{p_{2}}F_{n}^{-1}(u)\mathrm{d}u\bigg{)}-\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u$
$=$ $\mathbb{E}\big{(}\mathrm{Rem}(p_{2};F,F_{n})\big{)}-\mathbb{E}\big{(}\mathrm{Rem}(p_{1};F,F_{n})\big{)}.$
Proof.
Theorem 5.1 with $F_{n}$ instead of $G$ implies Corollary 5.2. ∎
Example 5.3 (SRS).
When $F_{n}$ is $F_{n,\textsc{srs}}$, condition (2.15) is satisfied. Furthermore, condition (2.16) is satisfied when $F\in\mathcal{T}_{\varepsilon}$ for some $\varepsilon>0$ (recall Example 2.2 for details). Hence, we conclude from Corollary 5.2 that when $F\in\mathcal{T}_{\varepsilon}$, the estimator $\int_{p_{1}}^{p_{2}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ of the middle-layer integral $\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u$, although being consistent by Example 5.2, has the bias
$\mathrm{Bias}_{n,\textsc{srs}}^{\textsc{middle}}(p_{1},p_{2})=\mathbb{E}\big{(}\mathrm{Rem}(p_{2};F,F_{n,\textsc{srs}})\big{)}-\mathbb{E}\big{(}\mathrm{Rem}(p_{1};F,F_{n,\textsc{srs}})\big{)}\in\mathbb{R},$
which might be any real number. Note that a sufficient condition for $F\in\mathcal{T}_{\varepsilon}$ for some $\varepsilon>0$ is the requirement $F\in\mathcal{L}_{\varepsilon}$, that is, $\mathbb{E}(|X|^{\varepsilon})<\infty$ for some $\varepsilon>0$.
Corollary 5.3 (asymptotic distribution).
Let $F$ be any cdf, and let $F_{1},F_{2},\ldots$ be any sequence of cdf’s satisfying condition (2.18) with $p\in\{p_{1},p_{2}\}$ and condition (2.19) with some normalizing constants $A_{n}\to\infty$ when $n\to\infty$. Then
$A_{n}\left(\int_{p_{1}}^{p_{2}}F_{n}^{-1}(u)\mathrm{d}u-\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u\right)=A_{n}\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}F(x)-F_{n}(x)\big{)}\mathrm{d}x+o_{\mathbb{P}}(1)$ (5.9)
and, therefore, the convergence-in-distribution statement
$A_{n}\left(\int_{p_{1}}^{p_{2}}F_{n}^{-1}(u)\mathrm{d}u-\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}_{\textsc{middle}}(p_{1},p_{2})$ (5.10)
holds if and only if
$A_{n}\left(\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F_{n}(x)\mathrm{d}x-\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F(x)\mathrm{d}x\right)\stackrel{{\scriptstyle d}}{{\to}}-\mathcal{L}_{\textsc{middle}}(p_{1},p_{2}),$ (5.11)
where $\mathcal{L}_{\textsc{middle}}(p_{1},p_{2})$ is a random variable determined by statement (5.11).
Proof.
Using Theorem 5.1 with $F_{n}$ instead of $G$, we only need to show that $A_{n}\mathrm{Rem}(p;F,F_{n})$ converges in probability to 00, which we showed in the proof of Corollary 2.3. This establishes equation (5.9), which implies the equivalence of statements (5.10) and (5.11), and concludes the proof of Corollary 5.3. ∎
Example 5.4 (SRS).
When $F_{n}$ is $F_{n,\textsc{srs}}$, condition (2.19) is a consequence of the Kolmogorov-Smirnov theorem. Since condition (2.19) is satisfied, condition (2.18) is satisfied as well, provided that, according to Lemma 2.1, the quantile function $F^{-1}$ is continuous at the points $p_{1}$ and $p_{2}$. Furthermore, since the integral $\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F_{n}(x)\mathrm{d}x$ is the arithmetic mean of $n$ independent copies of random variable (5.7), we conclude that the integral $\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F_{n}(x)\mathrm{d}x$ satisfies the central limit theorem. In summary, therefore, when the quantile function $F^{-1}$ is continuous at the points $p_{1}$ and $p_{2}$, we have
$\sqrt{n}\left(\int_{p_{1}}^{p_{2}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u-\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}(0,\sigma_{\textsc{middle}}^{2}(p_{1},p_{2}))$ (5.12)
where $\sigma_{\textsc{middle}}^{2}(p_{1},p_{2})$ is the variance of the random variable $Y_{\textsc{middle}}(p_{1},p_{2})$. The variance can be expressed by the formula (see Note A.1)
$\sigma_{\textsc{middle}}^{2}(p_{1},p_{2})=\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}F(x\wedge y)-F(x)F(y)\big{)}\mathrm{d}x\mathrm{d}y.$ (5.13)
Computation 5.1 (SRS).
To obtain a convenient computational formula for the empirical middle-layer integral $\int_{p_{1}}^{p_{2}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$, we can start with the equation
$\int_{p_{1}}^{p_{2}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u=\int_{0}^{p_{2}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u-\int_{0}^{p_{1}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ (5.14)
and then apply, for example, computational formula (4.15) for the two empirical lower-layer integrals on the right-hand side of equation (5.14). (All empirical integrals are finite, and so manipulations like those in equation (5.14) do not pose any technical issues.) In particular, we obtain the following equations
$\int_{p_{1}}^{p_{2}}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ $={1\over n}\sum_{i=\lceil np_{1}\rceil+1}^{\lceil np_{2}\rceil}X_{i:n}-\left({\lceil np_{2}\rceil\over n}-p_{2}\right)X_{\lceil np_{2}\rceil:n}+\left({\lceil np_{1}\rceil\over n}-p_{1}\right)X_{\lceil np_{1}\rceil:n}$
$={\lceil np_{2}\rceil-\lceil np_{1}\rceil\over n}\underbrace{{1\over\lceil np_{2}\rceil-\lceil np_{1}\rceil}\sum_{i=\lceil np_{1}\rceil+1}^{\lceil np_{2}\rceil}X_{i:n}}_{\textrm{trimmed mean}}$
$\qquad\underbrace{-\left({\lceil np_{2}\rceil\over n}-p_{2}\right)X_{\lceil np_{2}\rceil:n}+\left({\lceil np_{1}\rceil\over n}-p_{1}\right)X_{\lceil np_{1}\rceil:n}}_{\textrm{asymptotically negligible terms}}$
with the right-hand side connecting the empirical middle-layer integral with the trimmed (from both sides) mean considered by Stigler (1973).
6 Illustrations
To make the following illustrations less voluminous and easily connectable to what is known in the literature, and therefore maximally illuminating, we use SRS throughout this section. That is, we work with iid random variables $X_{1},\dots,X_{n}$ and from them arising empirical cdf $F_{n,\textsc{srs}}$, which is a consistent and unbiased estimator of the underlying population cdf $F$.
Example 6.1 (actuarial tail-value-at-risk).
Given any $p\in(0,1)$ and $F\in\mathcal{L}^{+}_{1}$, the actuarial tail-value-at-risk is
$\mathrm{TVaR}(p)={1\over 1-p}\int_{p}^{1}F^{-1}(u)\mathrm{d}u.$ (6.1)
As an illustration, in Figure 1(a) we depicted $p\mapsto\mathrm{TVaR}(p)$ for the Pareto (Type I) distribution, whose quantile function is $p\mapsto x_{0}(1-p)^{-1/\alpha}$ with the scale $x_{0}>0$ and shape $\alpha>0$ parameters, which we set to $1$ and $3$, respectively. The empirical counterpart $\mathrm{TVaR}_{n,\textsc{srs}}(p)$ is defined by equation (6.1) with $F_{n,\textsc{srs}}$ instead of $F$. (For a comprehensive review of various estimators of $\mathrm{TVaR}$ and its sister risk measure called Expected Shortfall, we refer to Nadarajah et al. (2013).) We have the following three statements concerning consistency, bias, and asymptotic normality of $\mathrm{TVaR}_{n,\textsc{srs}}(p)$.
Consistency.
When $F\in\mathcal{L}^{+}_{1}$, we obtain from Example 2.1 that $\mathrm{TVaR}_{n,\textsc{srs}}(p)$ is a consistent estimator of $\mathrm{TVaR}(p)$, that is,
$\mathrm{TVaR}_{n,\textsc{srs}}(p)\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\mathrm{TVaR}(p).$ (6.2)
This result is a special case of Gribkova et al. (2022a) who consider consistency and with it associated fixed-margin confidence intervals for the tail conditional allocation, which generalizes $\mathrm{TVaR}(p)$. For related bootstrap-type considerations and fixed-margin confidence intervals, we refer to Gribkova et al. (2024).
Bias.
When $F\in\mathcal{L}^{+}_{1}\cap\mathcal{T}_{\varepsilon}$ for some $\varepsilon>0$ (recall Note 2.3), we obtain from Example 2.2 that $\mathrm{TVaR}_{n,\textsc{srs}}(p)$ is a non-positively biased estimator of $\mathrm{TVaR}(p)$, that is,
$\mathbb{E}\big{(}\mathrm{TVaR}_{n,\textsc{srs}}(p)\big{)}\leq\mathrm{TVaR}(p)$ (6.3)
for every $n\in\mathbb{N}$. This property, although not stated explicitly, was earlier established by Brazauskas et al. (2008, p. 3602). García-Risueño (2025) offers an illuminating discussion of the property from the heuristic point of view.
Asymptotic normality.
When $F\in\mathcal{L}^{+}_{2}$ and the quantile function $F^{-1}$ is continuous at the point $p$, we obtain from Example 2.3 that
$\sqrt{n}\left(\mathrm{TVaR}_{n,\textsc{srs}}(p)-\mathrm{TVaR}(p)\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}(0,\sigma_{\mathrm{TVaR}}^{2}(p)),$ (6.4)
where
$\sigma_{\mathrm{TVaR}}^{2}(p)={1\over(1-p)^{2}}\int^{\infty}_{F^{-1}(p)}\int^{\infty}_{F^{-1}(p)}\big{(}F(x\wedge y)-F(x)F(y)\big{)}\mathrm{d}x\mathrm{d}y.$ (6.5)
This statement is a special case of a more general result derived by Brazauskas et al. (2008), who consider not just pointwise convergence to the limiting distribution but also uniform over all $p\in(0,1)$, which allows one to establish confidence bands for the function $p\mapsto\mathrm{TVaR}(p)$. Statement (6.4) is also a special case of Gribkova et al. (2022b) who consider the limiting distribution of the tail conditional allocation, which generalizes $\mathrm{TVaR}(p)$.
Example 6.2 (financial tail-value-at-risk).
Given $p\in(0,1)$ and $F\in\mathcal{L}^{-}_{1}$, the financial tail-value-at-risk is
$\mathrm{TV@R}(p)={1\over p}\int_{0}^{p}F^{-1}(u)\mathrm{d}u.$ (6.6)
As an illustration, in Figure 1(b) we depicted $p\mapsto\mathrm{TV@R}(p)$ for the Pareto (Type I) distribution with the scale $x_{0}=1$ and shape $\alpha=3$ parameters. The empirical counterpart $\mathrm{TV@R}_{n,\textsc{srs}}(p)$ is defined by equation (6.6) with $F_{n,\textsc{srs}}$ instead of $F$. We have the following three statements concerning consistency, bias, and asymptotic normality of $\mathrm{TV@R}_{n,\textsc{srs}}(p)$.
Consistency.
When $F\in\mathcal{L}^{-}_{1}$, we obtain from Example 4.1 that $\mathrm{TV@R}_{n,\textsc{srs}}(p)$ is a consistent estimator of $\mathrm{TV@R}(p)$, that is,
$\mathrm{TV@R}_{n,\textsc{srs}}(p)\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\mathrm{TV@R}(p).$ (6.7)
Bias.
When $F\in\mathcal{L}^{-}_{1}\cap\mathcal{T}_{\varepsilon}$ for some $\varepsilon>0$ (recall Note 4.3), we obtain from Example 4.2 that $\mathrm{TV@R}_{n,\textsc{srs}}(p)$ is a non-negatively biased estimator of $\mathrm{TV@R}(p)$, that is,
$\mathbb{E}\big{(}\mathrm{TV@R}_{n,\textsc{srs}}(p)\big{)}\geq\mathrm{TV@R}(p)$ (6.8)
for every $n\in\mathbb{N}$.
Asymptotic normality.
When $F\in\mathcal{L}^{-}_{2}$ and the quantile function $F^{-1}$ is continuous at the point $p$, we obtain from Example 4.3 that
$\sqrt{n}\left(\mathrm{TV@R}_{n,\textsc{srs}}(p)-\mathrm{TV@R}(p)\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}(0,\sigma_{\mathrm{TV@R}}^{2}(p)),$ (6.9)
where
$\sigma_{\mathrm{TV@R}}^{2}(p)={1\over p^{2}}\int_{-\infty}^{F^{-1}(p)}\int_{-\infty}^{F^{-1}(p)}\big{(}F(x\wedge y)-F(x)F(y)\big{)}\mathrm{d}x\mathrm{d}y.$ (6.10)
Example 6.3 (Lorenz curve).
Let $F\in\mathcal{L}^{\mu\neq 0}_{1}$, where $\mathcal{L}^{\mu\neq 0}_{1}$ denotes the set of all cdf’s that have finite but non-zero first moments, that is, $\mu=\mathbb{E}(X)\in\mathbb{R}\setminus\{0\}$ for $X\sim F$. The Lorenz curve (LC) is defined by (Gastwirth, 1971)
$\mathrm{LC}(p)={1\over\mu}\int_{0}^{p}F^{-1}(u)\mathrm{d}u,$ (6.11)
where $p$ runs through the unit interval $[0,1]$. (Usually, the Lorenz curve is defined and meaningfully interpreted only on the set of non-negative random variables $X\geq 0$.) Since the cases $p=0$ and $p=1$ are trivial, we restrict ourselves to $p\in(0,1)$. As an illustration, in Figure 1(c) we depicted $p\mapsto\mathrm{LC}(p)$ in the case of the Pareto (Type I) distribution with the scale $x_{0}=1$ and shape $\alpha=3$ parameters. When $F$ is replaced by $F_{n,\textsc{srs}}$ on the right-hand side of equation (6.11), we obtain the empirical Lorenz curve $\mathrm{LC}_{n,\textsc{srs}}(p)$. We have the following three statements concerning consistency, bias, and asymptotic normality of $\mathrm{LC}_{n,\textsc{srs}}(p)$ for any fixed $p\in(0,1)$.
Consistency.
When $F\in\mathcal{L}^{\mu\neq 0}_{1}$, we obtain from Example 4.1 and consistency of the arithmetic mean $\mu_{n,\textsc{srs}}$ that $\mathrm{LC}_{n,\textsc{srs}}(p)$ is a consistent estimator of $\mathrm{LC}(p)$, that is,
$\mathrm{LC}_{n,\textsc{srs}}(p)\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\mathrm{LC}(p).$ (6.12)
Bias.
The result of Example 4.2 suggests that $\mathrm{LC}_{n,\textsc{srs}}(p)$ might be a positively biased estimator of $\mathrm{LC}(p)$, given that the arithmetic mean $\mu_{n,\textsc{srs}}$ is an unbiased estimator of $\mu$. This may or may not be true as demonstrated by Arnold and Villaseñor (2015), who summarize their findings on the subject as follows:
For several parent distributions it is possible to prove that the sample Lorenz curve is a positively biased estimate of the population Lorenz curve. In this paper, several sufficient conditions for such positive bias are investigated. An example shows that negative bias is not impossible, though apparently not common. (Arnold and Villaseñor, 2015, p. 3)
Asymptotic normality.
When $F\in\mathcal{L}^{\mu\neq 0}_{2}$ and the quantile function $F^{-1}$ is continuous at the point $p$, we obtain from results of Section 4 and equation (1.1) that (see Lemma A.3 in Appendix A.3 for a proof)
$\sqrt{n}\left(\mathrm{LC}_{n,\textsc{srs}}(p)-\mathrm{LC}(p)\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}(0,\sigma_{\textsc{lc}}^{2}(p)),$ (6.13)
where $\sigma_{\textsc{lc}}^{2}(p)$ is the second moment of the mean-zero random variable
$Y_{\textsc{lc}}(p)={1\over\mu}\int_{-\infty}^{F^{-1}(p)}\big{(}\mathds{1}\{X\leq x\}-F(x)\big{)}\mathrm{d}x+{\mathrm{LC}(p)\over\mu}\big{(}X-\mu\big{)}.$ (6.14)
For confidence bands (i.e., simultaneous confidence intervals over all $p\in(0,1)$) for the Lorenz curve under minimal conditions on the population distribution, we refer to Csörgő et al. (1998). It is also useful to note that the variance $\sigma_{\textsc{lc}}^{2}(p)$ can be expressed in a form resembling those for $\sigma_{\mathrm{TVaR}}^{2}(p)$ and $\sigma_{\mathrm{TV@R}}^{2}(p)$ in equations (6.5) and (6.10), respectively, but apart from being nice theoretical exercises, we do not see much practical value in doing so. First, this would not lead to a convenient empirical estimator for $\sigma_{\textsc{lc}}^{2}(p)$. Second, the variance $\sigma_{\textsc{lc}}^{2}(p)$ can be easier estimated via a resampling technique (e.g., Shao and Tu, 1995).
Example 6.4 (Gini curve).
Let $F\in\mathcal{L}^{\mu\neq 0}_{1}$, where the class $\mathcal{L}^{\mu\neq 0}_{1}$ is defined in Example 6.3. The Gini curve (GC) is given by
$\mathrm{GC}(p)={1\over\mu}\left(\int_{1-p}^{1}F^{-1}(u)\mathrm{d}u-\int_{0}^{p}F^{-1}(u)\mathrm{d}u\right),$ (6.15)
where $p$ runs through the unit interval $[0,1]$. (Usually, the Gini curve is defined and meaningfully interpreted only on the set of non-negative random variables $X\geq 0$.) Since the cases $p=0$ and $p=1$ are trivial, we restrict ourselves to $p\in(0,1)$. As an illustration, in Figure 1(d) we depicted $p\mapsto\mathrm{GC}(p)$ in the case of the Pareto (Type I) distribution with the scale $x_{0}=1$ and shape $\alpha=3$ parameters. When $F$ is replaced by $F_{n,\textsc{srs}}$ on the right-hand side of equation (6.15), we obtain the empirical Gini curve $\mathrm{GC}_{n,\textsc{srs}}(p)$. We have the following three statements concerning consistency, bias, and asymptotic normality of $\mathrm{GC}_{n,\textsc{srs}}(p)$ for any $p\in(0,1)$.
Consistency.
When $F\in\mathcal{L}^{\mu\neq 0}_{1}$, we obtain from consistency of the arithmetic mean $\mu_{n,\textsc{srs}}$ and Examples 2.1 and 4.1 that $\mathrm{GC}_{n,\textsc{srs}}(p)$ is a consistent estimator of $\mathrm{GC}(p)$, that is,
$\mathrm{GC}_{n,\textsc{srs}}(p)\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\mathrm{GC}(p).$ (6.16)
Bias.
Even more so than in the case of the Lorenz curve, we cannot say anything definitive about the bias of $\mathrm{GC}_{n,\textsc{srs}}(p)$.
Asymptotic normality.
When $F\in\mathcal{L}^{\mu\neq 0}_{2}$ and the quantile function $F^{-1}$ is continuous at the points $p$ and $1-p$, we obtain from results of Section 4 and equation (1.1) that (see Lemma A.4 in Appendix A.3 for a proof)
$\sqrt{n}\left(\mathrm{GC}_{n,\textsc{srs}}(p)-\mathrm{GC}(p)\right)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}(0,\sigma_{\textsc{gc}}^{2}(p)),$ (6.17)
where $\sigma_{\textsc{gc}}^{2}(p)$ is the second moment of the mean-zero random variable
$Y_{\textsc{gc}}(p)={1\over\mu}\int_{-\infty}^{F^{-1}(1-p)}\big{(}\mathds{1}\{X\leq x\}-F(x)\big{)}\mathrm{d}x\\ +{1\over\mu}\int_{-\infty}^{F^{-1}(p)}\big{(}\mathds{1}\{X\leq x\}-F(x)\big{)}\mathrm{d}x+{1-\mathrm{GC}(p)\over\mu}\big{(}X-\mu\big{)}.$ (6.18)
7 $L$-functionals
Given what we have established in the previous sections, we may now wish to see if those results could be extended to the class of so-called $L$-functionals, whose definition we recall next. Namely, the $L$-functional $L_{w}:\mathcal{L}\to\overline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty,\infty\}$ is given by the equation
$L_{w}(F)=\int_{0}^{1}F^{-1}(u)w(u)\mathrm{d}u,$
where $w:(0,1)\to\mathbb{R}$ is a function. Next are several examples to illustrate what $w$’s arise in applications.
Example 7.1.
Let $\mathcal{G}\subset\mathcal{L}$ be the family of normal (i.e., Gaussian) distributions. To assess the scale parameter when location is known (e.g., Serfling, 1980, p. 269), the $L$-functional $L_{w}:\mathcal{G}\to\mathbb{R}$ is used with the weight function
$w(u)=\Phi^{-1}(u),$
where $\Phi$ denotes the standard normal cdf.
Example 7.2.
Let $\mathcal{L}\subset\mathcal{L}$ be the family of logistic distributions. To assess the location parameter when scale is known(e.g., Serfling, 1980, p. 269), the $L$-functional $L_{w}:\mathcal{L}\to\mathbb{R}$ is used with the weight function
$w(u)=6u(1-u).$
Example 7.3.
The Gini Mean Difference (GMD) is the $L$-functional $L_{w}:\mathcal{L}\to\overline{\mathbb{R}}$ with the weight function
$w(u)=4u-2.$
For properties and manifold applications of the GMD in economics and finance, we refer to Yitzhaki and Schechtman (2013).
Example 7.4.
The Tail Gini (Furman et al., 2017, p. 74) is the $L$-functional $L_{w}:\mathcal{L}\to\overline{\mathbb{R}}$ with the weight function
$w(u)=\frac{\mathds{1}\{p\leq u<1\}}{(1-p)^{2}}\big{(}4u-2(1+p)\big{)},$
where $p\in[0,1)$ is a parameter.
Example 7.5.
The Gini Shortfall (Furman et al., 2017, p. 75) is the $L$-functional $L_{w}:\mathcal{L}\to\overline{\mathbb{R}}$ with the weight function
$w(u)=\frac{\mathds{1}\{p\leq u<1\}}{(1-p)^{2}}\bigg{(}1-p+4\lambda\bigg{(}u-\frac{1+p}{2}\bigg{)}\bigg{)},$
where $p\in[0,1)$ and $\lambda\in[0,\infty)$ are parameters.
To give a flavour of the technical path that is often taken in the literature (e.g., Serfling, 1980, p. 265) when deriving large-sample statistical properties for $L$-integrals, we write the equations
$L_{w}(G)-L_{w}(F)$ $=\int_{0}^{1}\Big{(}G^{-1}(u)-F^{-1}(u)\Big{)}w(u)\mathrm{d}u$
$=\int_{\mathbb{R}}\Big{(}K_{w}\big{(}F(x)\big{)}-K_{w}\big{(}G(x)\big{)}\Big{)}\mathrm{d}x$ (7.1)
(see Appendix A.6.4 for technical details), where $K_{w}(t)=\int_{0}^{t}w(u)\mathrm{d}u$, assuming that $w$ is integrable on $(0,1)$. Note that equation (7.1) collapses into equation (1.1) when $w(u)\equiv 1$. With an estimator $F_{n}$ instead of $G$, we can now establish asymptotic properties of the empirical $L$-estimator $L_{w}(F_{n})$ when $n\to\infty$, but this path requires some form of the Taylor expansion applied on the function $K_{w}$, which imposes smoothness conditions on the weight function $w$. Although in some applications such conditions are satisfied, in many other applications they are not.
As a way out of the difficulty, we can successfully use our developed theory for integrated quantiles, and the path toward achieving this goal is analogous to that given by equation (7.1), although instead of integrating the weight function $w$ as in the definition of $K_{w}$, we now integrate the quantile function $F^{-1}$. The following example offers an idea how to achieve this goal.
Example 7.6 (a proof of concept).
For simplicity, let $F\in\mathcal{L}^{+}_{1}$, which means that integrals $\int_{p}^{1}F^{-1}(u)\mathrm{d}u$ are finite for every $p\in(0,1)$, and let the weight function $w$ be the difference $w_{1}-w_{2}$ of two non-decreasing, right-continuous, and non-negative functions $w_{1}$ and $w_{2}$. An example of such a weight function $w$ would be $w(u)=6u(1-u)$ that appears in Example 7.2. Hence,
$L_{w}(F)$ $=\int_{0}^{1}F^{-1}(u)w_{1}(u)\mathrm{d}u-\int_{0}^{1}F^{-1}(u)w_{2}(u)\mathrm{d}u$
$=\int_{0}^{\infty}\bigg{(}\int_{w_{1}^{-1}(x)}^{1}F^{-1}(u)\mathrm{d}u\bigg{)}\mathrm{d}x-\int_{0}^{\infty}\bigg{(}\int_{w_{2}^{-1}(x)}^{1}F^{-1}(u)\mathrm{d}u\bigg{)}\mathrm{d}x,$ (7.2)
with the technical details justifying the right-most equation given in Appendix A.6.5. Equation (7.2) paves a path that connects the results of Section 2 with asymptotic properties of the difference $L_{w}(F_{n})-L_{w}(F)$, where $F_{n}$, $n\in\mathbb{N}$, is a sequence of cdf’s that approximate $F$ when $n$ increases.
The following theorem, whose proof is given in Appendix A.4, implements the idea of Example 7.6 under minimal assumptions on $F$ and $w$, thus enabling to tackle $L$-functionals that arise from a whole spectrum of applications. To formulate the theorem, we use the notation
$L_{w,a,b}(F)=\int_{a}^{b}F^{-1}(u)w(u)\mathrm{d}u.$
Note that $L_{w}(F)=L_{w,a,b}(F)$ when $a=0$ and $b=1$.
Theorem 7.1.
Let the weight function $w$ be non-decreasing and right-continuous on the interval $(a,b)\subseteq(0,1)$. Then, for every cdf $F$ for which $L_{w,a,b}(F)$ is finite, we have the representation
$L_{w,a,b}(F)=-\int_{-\infty}^{0}\mathds{1}\{w(a)<x\}\Bigg{(}\int_{a}^{b\wedge w^{-1}(x)}F^{-1}(u)\mathrm{d}u\Bigg{)}\mathrm{d}x\\ +\int_{0}^{\infty}\mathds{1}\{x\leq w(b)\}\Bigg{(}\int_{a\vee w^{-1}(x)}^{b}F^{-1}(u)\mathrm{d}u\Bigg{)}\mathrm{d}x,$ (7.3)
where $w^{-1}$ is the left-continuous inverse of $w$.
We see from the right-hand side of equation (7.3) that when $a=0$ and $b=1$, the $L$-integral $L_{w,a,b}(F)$, which is equal to $L_{w}(F)$, can be expressed in terms of integrated quantiles under the assumption that the weight function $w$ is non-decreasing on the entire unit interval $(0,1)$, but this is a rather restrictive assumption in many applications. For this reason, we shall show next that all $w$’s of practical relevance can be expressed in terms of linear combinations of non-decreasing functions on various subintervals of $(0,1)$. This reduces $L_{w}(F)$ to a linear combinations of $L$-integrals $L_{w,a,b}(F)$ with various non-decreasing weight functions $w$ over various intervals $(a,b)\subset(0,1)$. To develop statistical inference for the $L$-integrals $L_{w,a,b}(F)$, we use the equation
$L_{w,a,b}(G)-L_{w,a,b}(F)=-\int_{-\infty}^{0}\mathds{1}\{w(a)<x\}\Bigg{(}\int_{a}^{b\wedge w^{-1}(x)}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u\Bigg{)}\mathrm{d}x\\ +\int_{0}^{\infty}\mathds{1}\{x\leq w(b)\}\Bigg{(}\int_{a\vee w^{-1}(x)}^{b}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u\Bigg{)}\mathrm{d}x$
that follows from Theorem 7.1. Using appropriate results from Sections 2, 4, and 5, we can now convert integrals of $G^{-1}-F^{-1}$ into integrals of $F-G$ for which statistical inference with $F_{n}$ instead of $G$ becomes straightforward.
Example 7.7.
Consider the weight function
$w(u)=6u(1-u),$
which we have encountered in Example 7.2. Obviously, we can decompose this weight function as the difference of $6u$ and $6u^{2}$ on the entire interval $(0,1)$. Alternatively, we can decompose it as the difference of $-6(1-u)^{2}$ and $-6(1-u)$ on the entire interval $(0,1)$. However, if we adopt the first decomposition, then we shall need to assume the finiteness of the moment $\mathbb{E}(X^{+})$, and if we adopt the second decomposition, then we shall need to assume the finiteness of the moment $\mathbb{E}(X^{-})$. Either way, superfluous conditions arise, which are not required for the finiteness of the $L$-integral $L_{w}(F)$ because the original weight function $w(u)$ converges to 00 when $u\downarrow 0$ and $u\uparrow 1$. For this reason, to reduce the problem to non-decreasing functions, we split the interval $[0,1)$ into two parts: $[0,1/2)$ where the function $w$ is increasing, and $[1/2,1)$ where the function $w$ is decreasing and thus $-w$ is increasing. Hence, we can express $w$ as the sum of two differences of non-decreasing functions, and this can be done in several ways. One way is to use the equation
$w(u)=\mathds{1}\{0\leq u<1/2\}\Big{(}w_{11}(u)-w_{12}(u)\Big{)}+\mathds{1}\{1/2\leq u<1\}\Big{(}w_{21}(u)-w_{22}(u)\Big{)}$ (7.4)
with the functions
$w_{11}(u)=w(u)\quad\textrm{and}\quad w_{12}(u)=0,$
$w_{21}(u)=0\quad\textrm{and}\quad w_{22}(u)=-w(u).$
Another way is to use equation (7.4) with the functions
$w_{11}(u)=6u\quad\textrm{and}\quad w_{12}(u)=6u^{2},$
$w_{21}(u)=-6(1-u)^{2}\quad\textrm{and}\quad w_{22}(u)=-6(1-u).$
Note that irrespective of the way we use, we always have the bounds
$|w_{11}(u)|\vee|w_{12}(u)|\leq c|w(u)|\quad\textrm{for all}\quad u\in[0,1/2),$
$|w_{21}(u)|\vee|w_{22}(u)|\leq c|w(u)|\quad\textrm{for all}\quad u\in[1/2,1),$
with some positive constant $c<\infty$. Therefore, no superfluous moment conditions arise when developing statistical inference for the $L$-integrals based on the four functions $w_{ij}$. This concludes Example 7.7.
Example 7.7 naturally leads to the following condition, which plays a fundamental role in connecting statistical inference for the $L$-integral $L_{w}(F)$ with that for integrated quantiles.
  1. (Cw)
    Let the weight function $w:(0,1)\to\mathbb{R}$ be such that, for some integer $K\geq 1$, there is a partition $10=a_{0}<a_{1}<\cdots<a_{K}=$ 1 of the unit interval $(0,1)$ and also pairs $(w_{11},w_{12}),\dots,(w_{K1},w_{K2})$ of non-decreasing and right-continuous on $(0,1)$ functions such that, for a constant $c<\infty$,
    $|w_{11}(u)|\vee|w_{12}(u)|\leq c|w(u)|\quad\textrm{in a neighbourhood of $0$},$ (7.5)
    $|w_{K1}(u)|\vee|w_{K2}(u)|\leq c|w(u)|\quad\textrm{in a neighbourhood of $1$},$ (7.6)
    and
    $w(u)=\sum_{k=1}^{K}\mathds{1}\{a_{k-1}\leq u<a_{k}\}\Big{(}w_{k1}(u)-w_{k2}(u)\Big{)}.$ (7.7)
To appreciate condition (Cw), which may look somewhat complex at first sight, we observe that if the function $w:(0,1)\to\mathbb{R}$ is piecewise monotonic, then the assumption is clearly satisfied. All the weight functions that we have encountered in the literature are piecewise monotonic. In the very special case when $w$ is monotonic (i.e., either non-decreasing or non-increasing) on the entire domain $(0,1)$ of its definition, condition (Cw) is satisfied with $K=1$. The following example illustrates the case $K=2$. Interestingly, we shall see from following Theorem 7.2, whose proof is constructive and given in Appendix A.5, that if condition (Cw) holds with $K\geq 3$, then the condition can be reduced to the case $K=2$.
Example 7.8.
Consider the weight function
$w(u)=(1-u)^{\beta}$
defined for all $u\in[0,1)$ and parameterized by $\beta\in\mathbb{R}$. The function can be expressed as the difference of two non-decreasing functions as follows:
$w(u)=\underbrace{\mathds{1}\{\beta\leq 0\}(1-u)^{\beta}}_{w_{1}(u)}-\underbrace{(-1)\mathds{1}\{\beta>0\}(1-u)^{\beta}}_{w_{2}(u)}.$
The two functions $w_{1}$ and $w_{2}$ are continuous and non-decreasing, and satisfy the bound
$|w_{1}(u)|\vee|w_{2}(u)|\leq|w(u)|$
for all $u\in(0,1)$. That is, for the weight function $w$ we do not need to split the interval $(0,1)$ into parts, and we can therefore set $K=1$. This concludes Example 7.8.
Theorem 7.2.
If condition (Cw) is satisfied with $K\geq 3$, then it is also satisfied with $K=2$.
Theorem 7.2, whose proof is given in Appendix A.5, says that there is a constant $a\in[0,1]$ and two pairs $(w_{11},w_{12})$ and $(w_{21},w_{22})$ of non-decreasing and right-continuous functions such that bounds (7.5) and (7.6) are satisfied and the representation
$w(u)=\mathds{1}\{0\leq u<a\}\Big{(}w_{11}(u)-w_{12}(u)\Big{)}+\mathds{1}\{a\leq u<1\}\Big{(}w_{21}(u)-w_{22}(u)\Big{)}$ (7.8)
holds. Earlier in Example 7.7, we showed that the case $K=2$ cannot be generally reduced to $K=1$. In other words, from the theoretical point of view, it is sufficient to consider only the cases $K=1$ and $K=2$. From the practical point of view, this minimality property could sometimes be time consuming to realize, but this is not an important issue: we can always work with any $K\geq 1$ that we can easily obtain.
8 Conclusion
We have developed a general statistical inference theory for integrals of quantiles. The required conditions are formulated in terms of general approximating sequences of cdf’s $F_{n}$, which are not attached to any specific sampling design or dependence structure. To maximally illuminate the developed theory, we have illustrated the obtained results using simple random sampling, due to everyone’s familiarity with this sampling design and basic statistical results. In particular, we have discussed consistency, bias, and asymptotic normality of various integrals of quantiles and their combinations that arise in risk and economic-inequality measurement: the actuarial and financial tail-values-at-risk ($\mathrm{TVaR}$ and $\mathrm{TV@R}$, respectively), and the Lorenz and Gini curves. To facilitate the appreciation of theoretical results, we have illustrated some of them numerically. Finally, we have linked the herein developed theory for integrated quantiles with topics such as $L$-integrals, distortion and spectral risk measures.
Appendix A Technicalities
We have subdivided this appendix into several major parts: Appendix A.1 contains auxiliary lemmas that we shall need in Appendix A.2 for proving Theorems 2.1, 4.1, and 5.1. In Appendix A.3 we shall derive asymptotic expansions for the empirical Lorenz and Gini curves that are needed for establishing their asymptotic normality stated in Examples 6.3 and 6.4. Appendices A.4 and A.5 contain proofs of Theorems 7.1 and 7.2, respectively. In Appendix A.6 we prove several cursory notes made earlier in the paper when introducing and discussing the main results.
A.1 Auxiliary lemmas
In this appendix we shall derive formulas for, and establish finiteness of, the first and second moments of the random variables $Y_{\textsc{upper}}(p)$ and $Y_{\textsc{lower}}(p)$. We shall need these results in Appendix A.2. The use of Fubini’s theorem and distributional equations $X\stackrel{{\scriptstyle d}}{{=}}F^{-1}(U)$ and $Y\stackrel{{\scriptstyle d}}{{=}}G^{-1}(U)$ will permeate our considerations, where $U$ is a uniform on $[0,1]$ random variable and =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution.
Lemma A.1.
If $F\in\mathcal{L}^{+}_{1}$, then the random variable $Y_{\textsc{upper}}(p)$ has a finite first moment, and the following equations hold:
$\mathbb{E}\big{(}Y_{\textsc{upper}}(p)\big{)}$ $=\int^{\infty}_{F^{-1}(p)}\big{(}1-F(x)\big{)}\mathrm{d}x$ (A.1)
$=\int_{p}^{1}F^{-1}(u)\mathrm{d}u-(1-p)F^{-1}(p).$ (A.2)
If $F\in\mathcal{L}^{+}_{2}$, then $Y_{\textsc{upper}}(p)$ has a finite second moment, and formula (2.24) holds for the variance of $Y_{\textsc{upper}}(p)$.
Proof.
We start with the equations
$Y_{\textsc{upper}}(p)$ $=\big{(}X-F^{-1}(p)\big{)}^{+}$ (A.3)
=$\stackrel{{\scriptstyle d}}{{=}}\big{(}F^{-1}(U)-F^{-1}(p)\big{)}^{+}.$ (A.4)
The first moment of the random variable on the right-hand side of equation (A.4) is finite whenever $F\in\mathcal{L}^{+}_{1}$. We can now use Fubini’s theorem to interchange the expectation and integration operations on the right-hand side of the equation
$\mathbb{E}\big{(}Y_{\textsc{upper}}(p)\big{)}=\mathbb{E}\bigg{(}\int_{F^{-1}(p)}^{\infty}\mathds{1}\{X>x\}\mathrm{d}x\bigg{)}$
and in this way arrive at equation (A.1) because $\mathbb{E}\big{(}\mathds{1}\{X>x\}\big{)}=1-F(x)$. To prove equation (A.2), we write
$\int_{F^{-1}(p)}^{\infty}\big{(}1-F(x))\mathrm{d}x$ $=\mathbb{E}\bigg{(}\int_{F^{-1}(p)}^{\infty}\mathds{1}\{X>x\}\mathrm{d}x\bigg{)}$
$=\mathbb{E}\Big{(}\big{(}F^{-1}(U)-F^{-1}(p)\big{)}^{+}\Big{)}$
$=\int_{0}^{1}\big{(}F^{-1}(s)-F^{-1}(p)\big{)}^{+}\mathrm{d}s$
$=\int_{p}^{1}\big{(}F^{-1}(s)-F^{-1}(p)\big{)}\mathrm{d}s$
$=\int_{p}^{1}F^{-1}(s)\mathrm{d}s-(1-p)F^{-1}(p).$
To prove the second half of the lemma, we first note that in view of equation (A.4), the random variable $Y_{\textsc{upper}}(p)$ has a finite second moment whenever $F\in\mathcal{L}^{+}_{2}$. To prove equation (2.24), we use Fubini’s theorem and have
$\mathrm{Var}\left(Y_{\textsc{upper}}(p)\right)$ $=\mathrm{Var}\left(\int^{\infty}_{F^{-1}(p)}\mathds{1}\{X>x\}\mathrm{d}x\right)$
$=\mathbb{E}\left(\bigg{(}\int^{\infty}_{F^{-1}(p)}\big{(}\mathds{1}\{X>x\}-(1-F(x))\big{)}\mathrm{d}x\bigg{)}^{2}\right)$
$=\int^{\infty}_{F^{-1}(p)}\int^{\infty}_{F^{-1}(p)}\mathbb{E}\Big{(}\big{(}\mathds{1}\{X\leq x\}-F(x)\big{)}\big{(}\mathds{1}\{X\leq y\}-F(y)\big{)}\Big{)}\mathrm{d}x\mathrm{d}y$
$=\int^{\infty}_{F^{-1}(p)}\int^{\infty}_{F^{-1}(p)}\big{(}F(x\wedge y)-F(x)F(y)\big{)}\mathrm{d}x\mathrm{d}y.$
This completes the proof of Lemma A.1. ∎
Lemma A.2.
If $F\in\mathcal{L}^{-}_{1}$, then the random variable $Y_{\textsc{lower}}(p)$ has a finite first moment, and the following equations hold:
$\mathbb{E}\big{(}Y_{\textsc{lower}}(p)\big{)}$ $=\int_{-\infty}^{F^{-1}(p)}F(x)\mathrm{d}x$ (A.5)
$=pF^{-1}(p)-\int_{0}^{p}F^{-1}(u)\mathrm{d}u.$ (A.6)
If $F\in\mathcal{L}^{-}_{2}$, then $Y_{\textsc{lower}}(p)$ has a finite second moment, and formula (4.14) holds for the variance of $Y_{\textsc{lower}}(p)$.
Proof.
We start with the equations
$Y_{\textsc{lower}}(p)$ $=\big{(}F^{-1}(p)-X\big{)}^{+}$ (A.7)
=$\stackrel{{\scriptstyle d}}{{=}}\big{(}F^{-1}(p)-F^{-1}(U)\big{)}^{+}.$ (A.8)
The first moment of the random variable on the right-hand side of equation (A.8) is finite whenever $F\in\mathcal{L}^{-}_{1}$. We can now use Fubini’s theorem to interchange the expectation and integration operations on the right-hand side of the equation
$\mathbb{E}\big{(}Y_{\textsc{lower}}(p)\big{)}=\mathbb{E}\bigg{(}\int_{-\infty}^{F^{-1}(p)}\mathds{1}\{X\leq x\}\mathrm{d}x\bigg{)}$
and in this way arrive at equation (A.5) because $\mathbb{E}\big{(}\mathds{1}\{X\leq x\}\big{)}=F(x)$. To prove equation (A.6), we write
$\int_{-\infty}^{F^{-1}(p)}F(x)\mathrm{d}x$ $=\mathbb{E}\bigg{(}\int_{-\infty}^{F^{-1}(p)}\mathds{1}\{X\leq x\}\mathrm{d}x\bigg{)}$
$=\mathbb{E}\Big{(}\big{(}F^{-1}(p)-F^{-1}(U)\big{)}^{+}\Big{)}$
$=\int_{0}^{1}\big{(}F^{-1}(p)-F^{-1}(u)\big{)}^{+}\mathrm{d}u$
$=\int_{0}^{p}\big{(}F^{-1}(p)-F^{-1}(u)\big{)}\mathrm{d}u$
$=pF^{-1}(p)-\int_{0}^{p}F^{-1}(u)\mathrm{d}u.$
To prove the second half of the lemma, we first note that in view of equation (A.8), the random variable $Y_{\textsc{lower}}(p)$ has a finite second moment whenever $F\in\mathcal{L}^{-}_{2}$. To prove equation (4.14), we use Fubini’s theorem and have
$\mathrm{Var}\left(Y_{\textsc{lower}}(p)\right)$ $=\mathrm{Var}\left(\int_{-\infty}^{F^{-1}(p)}\mathds{1}\{X\leq x\}\mathrm{d}x\right)$
$=\mathbb{E}\left(\bigg{(}\int_{-\infty}^{F^{-1}(p)}\big{(}\mathds{1}\{X\leq x\}-F(x)\big{)}\mathrm{d}x\bigg{)}^{2}\right)$
$=\int_{-\infty}^{F^{-1}(p)}\int_{-\infty}^{F^{-1}(p)}\mathbb{E}\Big{(}\big{(}\mathds{1}\{X\leq x\}-F(x)\big{)}\big{(}\mathds{1}\{X\leq y\}-F(y)\big{)}\Big{)}\mathrm{d}x\mathrm{d}y$
$=\int_{-\infty}^{F^{-1}(p)}\int_{-\infty}^{F^{-1}(p)}\big{(}F(x\wedge y)-F(x)F(y)\big{)}\mathrm{d}x\mathrm{d}y.$
This completes the proof of Lemma A.2. ∎
Note A.1.
The asymptotic variance $\sigma_{\textsc{middle}}^{2}(p_{1},p_{2})$ in statement (5.12) is the variance of the random variable $Y_{\textsc{middle}}(p_{1},p_{2})$ defined by equation (5.7). Since both $F^{-1}(p_{1})$ and $F^{-1}(p_{2})$ are finite, the random variable $Y_{\textsc{middle}}(p_{1},p_{2})$ is bounded and thus has finite moments of all degrees irrespective of the cdf $F$. Note the equations
$Y_{\textsc{middle}}(p_{1},p_{2})$ $=\int_{-\infty}^{F^{-1}(p_{2})}\mathds{1}\{X\leq x\}\mathrm{d}x-\int_{-\infty}^{F^{-1}(p_{1})}\mathds{1}\{X\leq x\}\mathrm{d}x$
$=\big{(}F^{-1}(p_{2})-X\big{)}^{+}-\big{(}F^{-1}(p_{1})-X\big{)}^{+}.$
This difference of the two positive parts, which may not have finite variances individually, always has a finite variance (as well as all the moments of higher degrees) irrespective of the cdf $F$, because $10<p_{1}<p_{2}<$ 1. Equation (5.13) follows immediately from
$\Big{(}Y_{\textsc{middle}}(p_{1},p_{2})-\mathbb{E}\big{(}Y_{\textsc{middle}}(p_{1},p_{2})\big{)}\Big{)}^{2}\\ =\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}\mathds{1}\{X\leq x\}-F(x)\big{)}\big{(}\mathds{1}\{X\leq y\}-F(y)\big{)}\mathrm{d}x\mathrm{d}y$
and Fubini’s theorem. This concludes Note A.1.
A.2 Proofs of Theorems 2.1, 4.1, and 5.1
The three theorems deal with arbitrary cdf’s $F$ and $G$ that, at most, are allowed to satisfy only moment conditions, with no further assumptions. It should be noted at the outset that we may find some portions of the following proofs look like simple consequences of integration-by-parts and change-of-variable formulas of calculus, but this would be mathematically inaccurate because, strictly speaking, the quantile function $F^{-1}$ is not an ordinary inverse of the cdf $F$, unless the latter is strictly increasing and continuous, which is not necessarily the case in the context of the present paper. Indeed, keeping in mind that our general results need to include, as a special case, the empirical cdf $F_{n,\textsc{srs}}$, which is a step-wise function and thus has flat regions as well as jumps, good care (recall Wacker (2023)) needs to be taken to maintain the largest possible class of cdf’s. Hence, only moment-type assumptions are required, as specified in the formulation of theorems.
Proof of Theorem 2.1.
By Lemma A.1, we have
$\int_{p}^{1}F^{-1}(u)\mathrm{d}u=(1-p)F^{-1}(p)+\int^{\infty}_{F^{-1}(p)}\big{(}1-F(x)\big{)}\mathrm{d}x.$ (A.9)
Since the same equation also holds with $G$ instead of $F$, taking the difference between the two equations gives
$\int_{p}^{1}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u=(1-p)\big{(}G^{-1}(p)-F^{-1}(p)\big{)}\\ +\int^{\infty}_{G^{-1}(p)}\big{(}1-G(x)\big{)}\mathrm{d}x-\int^{\infty}_{F^{-1}(p)}\big{(}1-F(x)\big{)}\mathrm{d}x.$ (A.10)
Next we express the penultimate integral on the right-hand side of equation (A.10) as follows:
$\int^{\infty}_{G^{-1}(p)}\big{(}1-G(x)\big{)}\mathrm{d}x=\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}1-G(x)\big{)}\mathrm{d}x+\int^{\infty}_{F^{-1}(p)}\big{(}1-G(x)\big{)}\mathrm{d}x,$ (A.11)
where the first integral on the right-hand side is equal to $-\int_{F^{-1}(p)}^{G^{-1}(p)}\big{(}1-G(x)\big{)}\mathrm{d}x$ when $F^{-1}(p)<G^{-1}(p)$. Combining equations (A.10) and (A.11), we arrive at equation (2.2) with the remainder term
$\mathrm{Rem}(p;F,G)=(1-p)\big{(}F^{-1}(p)-G^{-1}(p)\big{)}-\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}1-G(x)\big{)}\mathrm{d}x$
that can be concisely written as the integral on the right-hand side of equation (2.3).
To show that $\mathrm{Rem}(p;F,G)$ is non-negative, we proceed as follows. If $F^{-1}(p)\geq G^{-1}(p)$, then $x\geq G^{-1}(p)$, which is equivalent to $G(x)\geq p$, and so
$\mathrm{Rem}(p;F,G)=\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}G(x)-p\big{)}\mathrm{d}x\geq 0.$
If $F^{-1}(p)<G^{-1}(p)$, then $x<G^{-1}(p)$, which is equivalent to $G(x)<p$, and so
$\mathrm{Rem}(p;F,G)=-\int^{G^{-1}(p)}_{F^{-1}(p)}\big{(}G(x)-p\big{)}\mathrm{d}x\geq 0.$
Hence, we always have $\mathrm{Rem}(p;F,G)\geq 0$.
Furthermore,
$\mathrm{Rem}(p;F,G)$ $=\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}G(x)-F(x)\big{)}\mathrm{d}x+\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}F(x)-p\big{)}\mathrm{d}x$
$=\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}G(x)-F(x)\big{)}\mathrm{d}x-\mathrm{Rem}(p;G,F)$
$\leq\int^{F^{-1}(p)}_{G^{-1}(p)}\big{(}G(x)-F(x)\big{)}\mathrm{d}x,$
with the last inequality holding because $\mathrm{Rem}(p;G,F)\geq 0$. This establishes upper bound (2.4). Bound (2.5) follows immediately. ∎
Proof of Theorem 4.1.
By Lemma A.2, we have the equation
$\int_{0}^{p}F^{-1}(u)\mathrm{d}u=pF^{-1}(p)-\int_{-\infty}^{F^{-1}(p)}F(x)\mathrm{d}x.$ (A.12)
Since the same equation also holds with $G$ instead of $F$, taking the difference between the two equations gives
$\int_{p}^{1}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u=p\big{(}G^{-1}(p)-F^{-1}(p)\big{)}-\int_{-\infty}^{G^{-1}(p)}G(x)\mathrm{d}x+\int_{-\infty}^{F^{-1}(p)}F(x)\mathrm{d}x.$ (A.13)
Next we express the penultimate integral on the right-hand side of equation (A.13) as follows:
$\int_{-\infty}^{G^{-1}(p)}G(x)\mathrm{d}x=\int_{-\infty}^{F^{-1}(p)}G(x)\mathrm{d}x+\int_{F^{-1}(p)}^{G^{-1}(p)}G(x)\mathrm{d}x,$ (A.14)
where the second integral on the right-hand side is equal to $-\int^{F^{-1}(p)}_{G^{-1}(p)}G(x)\mathrm{d}x$ when $F^{-1}(p)>G^{-1}(p)$. Combining equations (A.13) and (A.14), we arrive at equation (4.2) with the remainder term
$\mathrm{Rem}(p;F,G)$ $=p\big{(}G^{-1}(p)-F^{-1}(p)\big{)}-\int_{F^{-1}(p)}^{G^{-1}(p)}G(x)\mathrm{d}x$
$=-p\big{(}F^{-1}(p)-G^{-1}(p)\big{)}+\int^{F^{-1}(p)}_{G^{-1}(p)}G(x)\mathrm{d}x,$
that can be concisely written as the integral on the right-hand side of equation (2.3). This completes the proof of Theorem 4.1. ∎
Proof of Theorem 5.1.
We start with the equation
$\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u=p_{2}F^{-1}(p_{2})-p_{1}F^{-1}(p_{1})-\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F(x)\mathrm{d}x,$ (A.15)
which can be verified in many different ways, and whose validity does not require any condition on the cdf $F$. For example, we can prove equation (A.15) as follows.
Since $F^{-1}(p_{1})$ and $F^{-1}(p_{2})$ are finite, the integral $\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F(x)\mathrm{d}x$ is finite. With $X$ denoting a random variable whose cdf is $F$, and also noting that $X$ is equal in distribution to $F^{-1}(U)$, where $U$ is a uniform on the interval $[0,1]$ random variable, we have the equations
$\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F(x)\mathrm{d}x$ $=\mathbb{E}\bigg{(}\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\mathds{1}\{X\leq x\}\mathrm{d}x\bigg{)}$
$=\mathbb{E}\Big{(}\big{(}F^{-1}(p_{2})-F^{-1}(p_{1})\vee X\big{)}^{+}\Big{)}$
$=\mathbb{E}\Big{(}\big{(}F^{-1}(p_{2})-F^{-1}(p_{1})\vee F^{-1}(U)\big{)}^{+}\Big{)}$
$=\int_{0}^{1}\big{(}F^{-1}(p_{2})-F^{-1}(p_{1}\vee u)\big{)}^{+}\mathrm{d}u.$
Splitting the integral $\int_{0}^{1}$ on the right-hand side into the sum of $\int_{0}^{p_{1}}$, $\int_{p_{1}}^{p_{2}}$ and $\int_{p_{2}}^{1}$, we obtain
$\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F(x)\mathrm{d}x$ $=\int_{0}^{p_{1}}\big{(}F^{-1}(p_{2})-F^{-1}(p_{1})\big{)}\mathrm{d}u+\int_{p_{1}}^{p_{2}}\big{(}F^{-1}(p_{2})-F^{-1}(u)\big{)}\mathrm{d}u$
$=p_{1}\big{(}F^{-1}(p_{2})-F^{-1}(p_{1})\big{)}+(p_{2}-p_{1})F^{-1}(p_{2})-\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u$
$=p_{2}F^{-1}(p_{2})-p_{1}F^{-1}(p_{1})-\int_{p_{1}}^{p_{2}}F^{-1}(u)\mathrm{d}u.$
Equation (A.15) follows. Of course, the same equation also holds with $G$ instead of $F$. Hence,
$\int_{p_{1}}^{p_{2}}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u=p_{2}\big{(}G^{-1}(p_{2})-F^{-1}(p_{2})\big{)}-p_{1}\big{(}G^{-1}(p_{1})-F^{-1}(p_{1})\big{)}\\ \quad-\bigg{(}\int_{G^{-1}(p_{1})}^{G^{-1}(p_{2})}G(x)\mathrm{d}x-\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}F(x)\mathrm{d}x\bigg{)},$
which can be rewritten as
$\int_{p_{1}}^{p_{2}}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u$ $=p_{2}\big{(}G^{-1}(p_{2})-F^{-1}(p_{2})\big{)}-p_{1}\big{(}G^{-1}(p_{1})-F^{-1}(p_{1})\big{)}$
$\quad-\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}G(x)-F(x)\big{)}\mathrm{d}x$
$\qquad-\bigg{(}\int_{G^{-1}(p_{1})}^{G^{-1}(p_{2})}G(x)\mathrm{d}x-\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}G(x)\mathrm{d}x\bigg{)}.$
Since
$\int_{G^{-1}(p_{1})}^{G^{-1}(p_{2})}G(x)\mathrm{d}x-\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}G(x)\mathrm{d}x=\int_{G^{-1}(p_{1})}^{F^{-1}(p_{1})}G(x)\mathrm{d}x-\int_{G^{-1}(p_{2})}^{F^{-1}(p_{2})}G(x)\mathrm{d}x,$
we have
$\int_{p_{1}}^{p_{2}}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u$ $=p_{2}\big{(}G^{-1}(p_{2})-F^{-1}(p_{2})\big{)}-p_{1}\big{(}G^{-1}(p_{1})-F^{-1}(p_{1})\big{)}$
$\quad-\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}G(x)-F(x)\big{)}\mathrm{d}x$
$\qquad-\int_{G^{-1}(p_{1})}^{F^{-1}(p_{1})}G(x)\mathrm{d}x+\int_{G^{-1}(p_{2})}^{F^{-1}(p_{2})}G(x)\mathrm{d}x.$
Combining the terms on the right-hand side of the latter equation, we obtain
$\int_{p_{1}}^{p_{2}}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u=-\int_{F^{-1}(p_{1})}^{F^{-1}(p_{2})}\big{(}G(x)-F(x)\big{)}\mathrm{d}x\\ +\int_{G^{-1}(p_{1})}^{F^{-1}(p_{1})}\big{(}p_{1}-G(x)\big{)}\mathrm{d}x-\int_{G^{-1}(p_{2})}^{F^{-1}(p_{2})}\big{(}p_{2}-G(x)\big{)}\mathrm{d}x.$
This establishes equation (5.2) and completes the proof of Theorem 5.1. ∎
A.3 Asymptotic results
Lemma A.3.
Let $F\in\mathcal{L}^{\mu\neq 0}_{2}$, and let the quantile function $F^{-1}$ be continuous at the point $p\in(0,1)$. Then
$\sqrt{n}\left(\mathrm{LC}_{n,\textsc{srs}}(p)-\mathrm{LC}(p)\right)=-{1\over\sqrt{n}}\sum_{i=1}^{n}Y_{i,\mathrm{LC}}(p)+o_{\mathbb{P}}(1),$ (A.16)
where
$Y_{i,\mathrm{LC}}(p)={1\over\mu}\int_{-\infty}^{F^{-1}(p)}\big{(}\mathds{1}\{X_{i}\leq x\}-F(x)\big{)}\mathrm{d}x+{\mathrm{LC}(p)\over\mu}\big{(}X_{i}-\mu\big{)}.$
Proof.
We start with the generic equation
${\theta_{n}\over\mu_{n}}-{\theta\over\mu}=\underbrace{{1\over\mu}(\theta_{n}-\theta)-{\theta\over\mu^{2}}(\mu_{n}-\mu)}_{\textrm{main term}}+\underbrace{{\theta\over\mu^{2}\mu_{n}}(\mu_{n}-\mu)^{2}-{1\over\mu\mu_{n}}(\theta_{n}-\theta)(\mu_{n}-\mu)}_{\textrm{remainder term}},$ (A.17)
which we next use with the following quantities:
  • $\mu$ is the mean of $F$,
  • $\theta$ is the lower-layer integral defined by equation (1.6),
  • $\mu_{n}$ is the arithmetic mean $\mu_{n,\textsc{srs}}$ given by equation (4.8),
  • $\theta_{n}$ is the estimator $\theta_{n,\textsc{srs}}=\int_{0}^{p}F_{n,\textsc{srs}}^{-1}(u)\mathrm{d}u$ of $\theta$.
By the classical central limit theorem, we have
$\sqrt{n}(\mu_{n,\textsc{srs}}-\mu)=O_{\mathbb{P}}(1).$
Statement (6.9) implies
$\sqrt{n}(\theta_{n,\textsc{srs}}-\theta)=O_{\mathbb{P}}(1).$
Consequently, the remainder term on the right-hand side of equation (A.17) is of the order $o_{\mathbb{P}}(1/\sqrt{n})$. Hence,
$\sqrt{n}\left(\mathrm{LC}_{n,\textsc{srs}}(p)-\mathrm{LC}(p)\right)$ $=\sqrt{n}\left({\theta_{n,\textsc{srs}}\over\mu_{n,\textsc{srs}}}-{\theta\over\mu}\right)$
$={1\over\mu}\sqrt{n}(\theta_{n,\textsc{srs}}-\theta)-{\theta\over\mu^{2}}\sqrt{n}(\mu_{n,\textsc{srs}}-\mu)+o_{\mathbb{P}}(1).$ (A.18)
By equation (4.10) with $F_{n,\textsc{srs}}$ instead of $F_{n}$, we have
$\sqrt{n}(\theta_{n,\textsc{srs}}-\theta)=\sqrt{n}\,\int_{-\infty}^{F^{-1}(p)}\big{(}F(x)-F_{n,\textsc{srs}}(x)\big{)}\mathrm{d}x+o_{\mathbb{P}}(1).$ (A.19)
Combining equations (A.18) and (A.19), we obtain
$\sqrt{n}\left(\mathrm{LC}_{n,\textsc{srs}}(p)-\mathrm{LC}(p)\right)=-{1\over\mu}\sqrt{n}\,\int_{-\infty}^{F^{-1}(p)}\big{(}F_{n,\textsc{srs}}(x)-F(x)\big{)}\mathrm{d}x-{\theta\over\mu^{2}}\sqrt{n}(\mu_{n,\textsc{srs}}-\mu)+o_{\mathbb{P}}(1),$
which implies statement (A.16) and concludes the proof of Lemma A.3. ∎
Lemma A.4.
Let $F\in\mathcal{L}^{\mu\neq 0}_{2}$, and let the quantile function $F^{-1}$ be continuous at the points $p\in(0,1)$ and $1-p$. Then
$\sqrt{n}\left(\mathrm{GC}_{n,\textsc{srs}}(p)-\mathrm{GC}(p)\right)={1\over\sqrt{n}}\sum_{i=1}^{n}\Big{(}Y_{i,\mathrm{LC}}(1-p)+Y_{i,\mathrm{LC}}(p)\Big{)}+o_{\mathbb{P}}(1),$
where
$Y_{i,\mathrm{LC}}(u)={1\over\mu}\int_{-\infty}^{F^{-1}(u)}\big{(}\mathds{1}\{X_{i}\leq x\}-F(x)\big{)}\mathrm{d}x+{\mathrm{LC}(u)\over\mu}\big{(}X_{i}-\mu\big{)}.$
Proof.
Note that
$\mathrm{GC}(p)$ $=1-{1\over\mu}\left(\int_{0}^{1-p}F^{-1}(u)\mathrm{d}u+\int_{0}^{p}F^{-1}(u)\mathrm{d}u\right).$
$=1-\mathrm{LC}(1-p)-\mathrm{LC}(p).$
Hence,
$\sqrt{n}\left(\mathrm{GC}_{n,\textsc{srs}}(p)-\mathrm{GC}(p)\right)=-\sqrt{n}\left(\mathrm{LC}_{n,\textsc{srs}}(1-p)-\mathrm{LC}(1-p)\right)-\sqrt{n}\left(\mathrm{LC}_{n,\textsc{srs}}(p)-\mathrm{LC}(p)\right),$
and with the help of Lemma A.3, we obtain
$\sqrt{n}\left(\mathrm{GC}_{n,\textsc{srs}}(p)-\mathrm{GC}(p)\right)={1\over\sqrt{n}}\sum_{i=1}^{n}\Big{(}Y_{i,\mathrm{LC}}(1-p)+Y_{i,\mathrm{LC}}(p)\Big{)}+o_{\mathbb{P}}(1),$
which concludes the proof of Lemma A.4. ∎
A.4 Proof of Theorem 7.1
We begin the proof by splitting the integral $L_{w,a,b}(F)$ into two parts according to whether $w(u)<0$ or $w(u)\geq 0$. This gives us the equation
$\int_{a}^{b}F^{-1}(u)w(u)\mathrm{d}u=\int_{a}^{b}F^{-1}(u)\mathds{1}\{w(u)<0\}w(u)\mathrm{d}u+\int_{a}^{b}F^{-1}(u)\mathds{1}\{w(u)\geq 0\}w(u)\mathrm{d}u.$ (A.20)
We shall next analyze the two integrals on the right-hand side separately, showing that they are equal to the respective integrals on the right-hand side of equation (7.3).
Pertaining to the first integral on the right-hand side of equation (A.20), since $w$ is non-decreasing and right-continuous, we have $w(u)<0$ if and only if $u<w^{-1}(0)$, and so
$\int_{a}^{b}F^{-1}(u)\mathds{1}\{w(u)<0\}w(u)\mathrm{d}u=\int_{a}^{b}F^{-1}(u)\mathds{1}\{u<w^{-1}(0)\}w(u)\mathrm{d}u.$
Furthermore, since $w(u)<0$, we have
$w(u)$ $=-\int_{w(u)}^{0}\mathrm{d}x$
$=-\int_{-\infty}^{0}\mathds{1}\{u<w^{-1}(x)\}\mathrm{d}x,$
where the last equation holds because we have $w(u)<x$ if and only if $u<w^{-1}(x)$, due to our assumption that $w$ is non-decreasing and right-continuous. Consequently,
$\int_{a}^{b}F^{-1}(u)\mathds{1}\{w(u)<0\}w(u)\mathrm{d}u$ $=-\int_{-\infty}^{0}\Bigg{(}\int_{a}^{b}F^{-1}(u)\mathds{1}\{u<w^{-1}(0)\wedge w^{-1}(x)\}\mathrm{d}u\Bigg{)}\mathrm{d}x$
$=-\int_{-\infty}^{0}\Bigg{(}\int_{a}^{b}F^{-1}(u)\mathds{1}\{u<w^{-1}(x)\}\mathrm{d}u\Bigg{)}\mathrm{d}x,$ (A.21)
where right-most equation holds because $w^{-1}$ is non-decreasing. The double integral on the right-hand side of equation (A.21) is zero when $w^{-1}(x)\leq a$. Hence, continuing with equation (A.21) for only those $x$ that satisfy $a<w^{-1}(x)$, we have
$\int_{a}^{b}F^{-1}(u)\mathds{1}\{w(u)<0\}w(u)\mathrm{d}u$ $=-\int_{-\infty}^{0}\mathds{1}\{a<w^{-1}(x)\}\Bigg{(}\int_{a}^{b}F^{-1}(u)\mathds{1}\{u<w^{-1}(x)\}\mathrm{d}u\Bigg{)}\mathrm{d}x$
$=-\int_{-\infty}^{0}\mathds{1}\{a<w^{-1}(x)\}\Bigg{(}\int_{a}^{b\wedge w^{-1}(x)}F^{-1}(u)\mathrm{d}u\Bigg{)}\mathrm{d}x.$ (A.22)
Since $w$ is non-decreasing and right-continuous, we have that $a<w^{-1}(x)$ is equivalent to $w(a)<x$, and so equation (A.22) gives the first integral on the right-hand side of equation (7.3).
Analogously we tackle the second integral on the right-hand side of equation (A.20). First, since $w(u)\geq 0$ if and only if $u\geq w^{-1}(0)$, we have
$\int_{a}^{b}F^{-1}(u)\mathds{1}\{w(u)\geq 0\}w(u)\mathrm{d}u=\int_{a}^{b}F^{-1}(u)\mathds{1}\{u\geq w^{-1}(0)\}w(u)\mathrm{d}u.$
Furthermore, since $w(u)\geq 0$, we have
$w(u)$ $=\int_{0}^{w(u)}\mathrm{d}x$
$=\int_{0}^{\infty}\mathds{1}\{u\geq w^{-1}(x)\}\mathrm{d}x,$
where the last equation holds because $w(u)\geq x$ if and only if $u\geq w^{-1}(x)$. Consequently,
$\int_{a}^{b}F^{-1}(u)\mathds{1}\{w(u)\geq 0\}w(u)\mathrm{d}u$ $=\int_{0}^{\infty}\Bigg{(}\int_{a}^{b}F^{-1}(u)\mathds{1}\{u\geq w^{-1}(0)\vee w^{-1}(x)\}\mathrm{d}u\Bigg{)}\mathrm{d}x$
$=\int_{0}^{\infty}\Bigg{(}\int_{a}^{b}F^{-1}(u)\mathds{1}\{u\geq w^{-1}(x)\}\mathrm{d}u\Bigg{)}\mathrm{d}x,$ (A.23)
where right-most equation holds because $w^{-1}$ is non-decreasing. The double integral on the right-hand side of equation (A.23) is zero when $w^{-1}(x)>b$. Hence, continuing with equation (A.23) for only those $x$ that satisfy $w^{-1}(x)\leq b$, we have
$\int_{a}^{b}F^{-1}(u)\mathds{1}\{w(u)\geq 0\}w(u)\mathrm{d}u$ $=\int_{0}^{\infty}\mathds{1}\{w^{-1}(x)\leq b\}\Bigg{(}\int_{a}^{b}F^{-1}(u)\mathds{1}\{u\geq w^{-1}(x)\}\mathrm{d}u\Bigg{)}\mathrm{d}x$
$=\int_{0}^{\infty}\mathds{1}\{w^{-1}(x)\leq b\}\Bigg{(}\int_{a\vee w^{-1}(x)}^{b}F^{-1}(u)\mathrm{d}u\Bigg{)}\mathrm{d}x.$ (A.24)
Since $w$ is non-decreasing and right-continuous, we have that $w^{-1}(x)\leq b$ is equivalent to $x\leq w(b)$, and so equation (A.24) gives the second integral on the right-hand side of equation (7.3). This concludes the proof of Theorem 7.1.
A.5 Proof of Theorem 7.2.
Let condition (Cw) be satisfied with some $K\geq 3$. Equation (7.7) is equivalent to
$w(u)$ $=\mathds{1}\{0\leq t<a_{1}\}\Big{(}w_{11}(u)-w_{12}(u)\Big{)}$
$\qquad+\sum_{k=2}^{K-1}\mathds{1}\{a_{k-1}\leq t<a_{k}\}\Bigg{(}\bigg{\{}w_{k1}(u)+\sum_{i=2}^{k}c_{i}\bigg{\}}-\bigg{\{}w_{k2}(u)+\sum_{i=2}^{k}c_{i}\bigg{\}}\Bigg{)}$
$\qquad+\mathds{1}\{a_{K-1}\leq t<1\}\Big{(}w_{K1}(u)-w_{K2}(u)\Big{)},$
where the constants $c_{2},\dots,c_{K-1}$ can be any real numbers, but we choose them so that the inequalities
$w_{k1}(a_{k-1})+\sum_{i=2}^{k}c_{i}\geq w_{(k-1)1}(a_{k-1}-)+\sum_{i=2}^{k-1}c_{i}\qquad\Big{(}\Longleftrightarrow w_{k1}(a_{k-1})+c_{k}\geq w_{(k-1)1}(a_{k-1}-)\Big{)}$
and
$w_{k2}(a_{k-1})+\sum_{i=2}^{k}c_{i}\geq w_{(k-1)2}(a_{k-1}-)+\sum_{i=2}^{k-1}c_{i}\qquad\Big{(}\Longleftrightarrow w_{k2}(a_{k-1})+c_{k}\geq w_{(k-1)2}(a_{k-1}-)\Big{)}$
would hold for every $k=2,\dots,K-1$. The functions
$w_{11}^{*}(u)=\mathds{1}\{0\leq u<a_{1}\}w_{11}(u)+\sum_{k=2}^{K-1}\mathds{1}\{a_{k-1}\leq u<a_{k}\}\bigg{\{}w_{k1}(u)+\sum_{i=2}^{k}c_{i}\bigg{\}}$
and
$w_{12}^{*}(u)=\mathds{1}\{0\leq u<a_{1}\}w_{12}(u)+\sum_{k=2}^{K-1}\mathds{1}\{a_{k-1}\leq u<a_{k}\}\bigg{\{}w_{k2}(u)+\sum_{i=2}^{k}c_{i}\bigg{\}}$
are non-decreasing and right-continuous on the interval $[0,a_{K-1})$. Furthermore, the functions
$w_{21}^{*}(u)=w_{K1}(u)\quad\textrm{and}\quad w_{22}^{*}(u)=w_{K2}(u)$
are non-decreasing and right-continuous on the interval $[a_{K-1},1)$. We obviously have the equation
$w(u)=\mathds{1}\{0\leq u<a_{K-1}\}\Big{(}w_{11}^{*}(u)-w_{12}^{*}(u)\Big{)}+\mathds{1}\{a_{K-1}\leq u<1\}\Big{(}w_{21}^{*}(u)-w_{22}^{*}(u)\Big{)},$
which establishes representation (7.8) and completes the proof of Theorem 7.2.
A.6 Proofs of cursory notes
In this appendix we present proofs of some of the cursory notes that we have made while discussing main results.
A.6.1 Proof of equation (1.1)
Let $U$ denote a uniform on $[0,1]$ random variable. Since $X$ and $Y$ have the same distributions as $F^{-1}(U)$ and $G^{-1}(U)$, respectively, we therefore have
$\int_{0}^{1}\big{(}G^{-1}(u)-F^{-1}(u)\big{)}\mathrm{d}u$ $=\mathbb{E}\big{(}G^{-1}(U)\big{)}-\mathbb{E}\big{(}F^{-1}(U)\big{)}$
$=\mathbb{E}(Y)-\mathbb{E}(X).$ (A.25)
Furthermore, using Fubini’s theorem, we have
$\int^{\infty}_{-\infty}\big{(}F(x)-G(x)\big{)}\mathrm{d}x$ $=\mathbb{E}\left(\int^{\infty}_{-\infty}\big{(}\mathds{1}\{X\leq x\}-\mathds{1}\{Y\leq x\}\big{)}\mathrm{d}x\right)$
$=\mathbb{E}\left(\int^{\infty}_{-\infty}\big{(}\mathds{1}\{Y>x\}-\mathds{1}\{X>x\}\big{)}\mathrm{d}x\right)$
$=\mathbb{E}(Y-X),$ (A.26)
where the right-most equation holds because the integral inside the expectation is equal to $Y-X$. Since the right-hand sides of equations (A.25) and (A.26) are equal, we have equation (1.1).
A.6.2 On the validity of statements (3.4) and (3.7)-(3.8)
Using the definition of the empirical quantile function $H_{n,\textrm{SRS}}^{-1}$, statement (3.4) is equivalent to
$Z_{\lceil n/2\rceil:n}\stackrel{{\scriptstyle\mathbb{P}}}{{\nrightarrow}}\frac{1}{2}.$
To prove it, we want to find $\epsilon>0$ such that
$\lim_{n\to\infty}\mathbb{P}\left(\left|Z_{\lceil n/2\rceil:n}-\frac{1}{2}\right|>\epsilon\right)>0.$ (A.27)
We set $\epsilon=1/2$ and note that statement (A.27) follows if we establish at least one of statements (3.7) and (3.8), but for completeness of the argument, we shall next establish both of them. Using a well-known formula (e.g., David and Nagaraja, 2003, eq. (2.1.3)) for the cdf of order statistics, we have
$\mathbb{P}\left(Z_{\lceil n/2\rceil:n}>\frac{3}{2}\right)$ $=1-\sum_{j=\lceil n/2\rceil}^{n}\mathbb{P}\left(Z_{j:n}\leq\frac{3}{2}\right)$
$=1-\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}\left(F_{Z}\left(\frac{3}{2}\right)\right)^{j}\left(1-F_{Z}\left(\frac{3}{2}\right)\right)^{n-j}$
$=1-\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}\left(\frac{1}{2}\right)^{j}\left(\frac{1}{2}\right)^{n-j}$
$=1-\left(\frac{1}{2}\right)^{n}\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}.$ (A.28)
Similarly,
$\mathbb{P}\left(Z_{\lceil n/2\rceil:n}<\frac{1}{2}\right)$ $=\sum_{j=\lceil n/2\rceil}^{n}\mathbb{P}\left(Z_{j:n}<\frac{1}{2}\right)$
$=\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}\left(F_{Z}\left(\frac{1}{2}\right)\right)^{j}\left(1-F_{Z}\left(\frac{1}{2}\right)\right)^{n-j}$
$=\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}\left(\frac{1}{2}\right)^{j}\left(\frac{1}{2}\right)^{n-j}$
$=\left(\frac{1}{2}\right)^{n}\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}.$ (A.29)
We next rearrange the sums of binomials on the right-hand sides of equations (A.28) and (A.29). When $n$ is even, $\lceil n/2\rceil=n/2$ and $2^{n}$ can be expressed in $n+1$ (which is odd) terms
$2^{n}$ $=\binom{n}{0}+\binom{n}{1}+\cdots+\binom{n}{\frac{n}{2}-1}+\binom{n}{\frac{n}{2}}+\binom{n}{\frac{n}{2}+1}+\cdots+\binom{n}{n-1}+\binom{n}{n}$
$=\binom{n}{\frac{n}{2}}+2\left(\binom{n}{\frac{n}{2}+1}+\cdots+\binom{n}{n-1}+\binom{n}{n}\right).$
Hence, we have
$\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}$ $=\binom{n}{\lceil\frac{n}{2}\rceil}+\binom{n}{\lceil\frac{n}{2}\rceil+1}+\cdots+\binom{n}{n-1}+\binom{n}{n}$
$=\binom{n}{\frac{n}{2}}+\binom{n}{\frac{n}{2}+1}+\cdots+\binom{n}{n-1}+\binom{n}{n}$
$={\frac{1}{2}}\left(2^{n}+\binom{n}{\frac{n}{2}}\right)$
$=2^{n-1}+{\frac{1}{2}}\binom{n}{\frac{n}{2}}.$
When $n$ is odd, $2^{n}$ can be expressed in $n+1$ (which is even) terms as follows:
$2^{n}$ $=\binom{n}{0}+\binom{n}{1}+\cdots+\binom{n}{\lceil\frac{n}{2}\rceil-1}+\binom{n}{\lceil\frac{n}{2}\rceil}+\cdots+\binom{n}{n-1}+\binom{n}{n}$
$=2\left(\binom{n}{\lceil\frac{n}{2}\rceil}+\cdots+\binom{n}{n-1}+\binom{n}{n}\right).$
Hence, we have
$\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}$ $=\binom{n}{\lceil\frac{n}{2}\rceil}+\binom{n}{\lceil\frac{n}{2}\rceil+1}+\cdots+\binom{n}{n-1}+\binom{n}{n}$
$={\frac{1}{2}}2^{n}$
$=2^{n-1}.$
Going back to the original expressions (A.28) and (A.29) of probabilities, we have
$\mathbb{P}\left(Z_{\lceil n/2\rceil:n}>\frac{3}{2}\right)$ $=1-\left(\frac{1}{2}\right)^{n}\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}$
=$=\begin{cases}\frac{1}{2}-\frac{1}{2^{n+1}}\binom{n}{n/2}&\textrm{when}\;n\;\textrm{is even};\\ \frac{1}{2}&\textrm{when}\;n\;\textrm{is odd};\end{cases}$
and
$\mathbb{P}\left(Z_{\lceil n/2\rceil:n}<\frac{1}{2}\right)$ $=\left(\frac{1}{2}\right)^{n}\sum_{j=\lceil n/2\rceil}^{n}\binom{n}{j}$
=$=\begin{cases}\frac{1}{2}+\frac{1}{2^{n+1}}\binom{n}{n/2}&\textrm{when}\;n\;\textrm{is even};\\ \frac{1}{2}&\textrm{when}\;n\;\textrm{is odd}.\end{cases}$
As $n\to\infty$, the term $\frac{1}{2^{n+1}}\binom{n}{n/2}$ vanishes. This establishes statements (3.7) and (3.8), and concludes the proof of statement (3.4).
A.6.3 Derivation of expression (3.11)
Starting with equation (3.10), we need to calculate the first two moments of the random variable $\big{(}H^{-1}(U)-(1-a)\big{)}^{+}$. From Figure 3.2 we see that the random variable $H^{-1}(U)-(1-a)$ is (strictly) positive if and only if $U>1/2$, and under this condition, $H^{-1}(U)$ is equal to $2(1-a)U+2a$. Hence,
$\big{(}H^{-1}(U)-(1-a)\big{)}^{+}=\big{(}2(1-a)U+2a\big{)}\mathds{1}\{U>1/2\}.$ (A.30)
By calculating the first two moments of the random variable on the right-hand side of equation (A.30), we arrive at expression (3.11).
A.6.4 Proof of equation (7.1)
We start with the equations
$K_{w}(F(x))$ $=\int_{0}^{F(x)}w(u)\mathrm{d}u$
$=\int_{0}^{1}\mathds{1}\{u\leq F(x)\}w(u)\mathrm{d}u$
$=\int_{0}^{1}\mathds{1}\{F^{-1}(u)\leq x\}w(u)\mathrm{d}u.$
Analogous equations hold for $K_{w}(G(x))$. Hence,
$\int_{\mathbb{R}}\Big{(}K_{w}\big{(}F(x)\big{)}-K_{w}\big{(}G(x)\big{)}\Big{)}\mathrm{d}x$ $=\int_{\mathbb{R}}\bigg{(}\int_{0}^{1}\Big{(}\mathds{1}\{F^{-1}(u)\leq x\}-\mathds{1}\{G^{-1}(u)\leq x\}\Big{)}w(u)\mathrm{d}u\bigg{)}\mathrm{d}x$
$=\int_{0}^{1}\bigg{(}\int_{\mathbb{R}}\Big{(}\mathds{1}\{F^{-1}(u)\leq x\}-\mathds{1}\{G^{-1}(u)\leq x\}\Big{)}\mathrm{d}x\bigg{)}w(u)\mathrm{d}u$
$=\int_{0}^{1}\Big{(}G^{-1}(u)-F^{-1}(u)\Big{)}w(u)\mathrm{d}u,$
where the right-most equation holds because the inner integral is equal to $G^{-1}(u)-F^{-1}(u)$. This completes the proof of equation (7.1).
A.6.5 Proof of equation (7.2)
Since $w_{1}$ is non-decreasing and right-continuous, and such that $w_{1}(0)\geq 0$, we have
$\int_{0}^{1}F^{-1}(u)w_{1}(u)\mathrm{d}u$ $=\int_{0}^{1}F^{-1}(u)\bigg{(}\int_{0}^{\infty}\mathds{1}\{x\leq w_{1}(u)\}\mathrm{d}x\bigg{)}\mathrm{d}u$
$=\int_{0}^{1}F^{-1}(u)\bigg{(}\int_{0}^{\infty}\mathds{1}\{w_{1}^{-1}(x)\leq u\}\mathrm{d}x\bigg{)}\mathrm{d}u$
$=\int_{0}^{\infty}\bigg{(}\int_{0}^{1}F^{-1}(u)\mathds{1}\{w_{1}^{-1}(x)\leq u\}\mathrm{d}u\bigg{)}\mathrm{d}x$
$=\int_{0}^{\infty}\bigg{(}\int_{w_{1}^{-1}(x)}^{1}F^{-1}(u)\mathrm{d}u\bigg{)}\mathrm{d}x.$
An analogous equations hold for $w_{2}$. This establishes equation (7.2).